### v1

# Can Open-source LLMs Enhance Data Augmentation for Toxic Detection?: An Experimental Study

Abstract
High-quality, diverse harmful data is essential to addressing real-time applications in content moderation. Current state-of-the-art approaches to toxic content detection using GPT series models are costly and lack explainability. This paper investigates the use of prompt engineering and fine-tuning techniques on open-source LLMs to enhance harmful data augmentation specifically for toxic content detection. We conduct a two-stage empirical study, with stage 1 evaluating six open-source LLMs across multiple datasets using only prompt engineering and stage 2 focusing on fine-tuning. Our findings indicate that Mistral can excel in generating harmful data with minimal hallucination. While fine-tuning these models improves data quality and diversity, challenges such as data duplication and overfitting persist. Our experimental results highlight scalable, cost-effective strategies for enhancing toxic content detection systems. These findings not only demonstrate the potential of open-source LLMs in creating robust content moderation tools. The application of this method in real industrial scenarios further proves the feasibility and efficiency of the fine-tuned open-source LLMs for data augmentation. We hope our study will aid in understanding the capabilities and limitations of current models in toxic content detection and drive further advancements in this field.
Warning: This paper has instances of hateful and offensive language to serve as examples.
1 Introduction
With the development of large language models Brown et al. (2020), there is a significant amount of work currently using LLMs for data augmentation Whitehouse et al. (2023); Ye et al. (2022); Wang et al. (2023). This study investigates the application of prompt engineering and fine-tuning techniques on open-source large language models to enhance the augmentation of harmful data, crucial for improving toxic content detection systems. Data augmentation for toxic detection is vital because real-world datasets often lack the diversity and volume needed to train robust models Garg et al. (2023). Current challenges in toxic detection include the need for real-time processing Oikawa et al. (2022), where toxic comments must be filtered out swiftly to prevent user exposure. Existing methods using LLMs to directly detect toxic content, while effective, often suffer from slow inference speeds and high computational resource requirements, making them unsuitable for systems without GPU resources Kumarage et al. (2024). Additionally, methods based on the GPT Brown et al. (2020) series, despite their good performance, face issues of high cost and lack of explainability Zhao et al. (2024). Content safety classifiers also often face urgency and specificity across various scenarios due to differing definitions of sensitive topics among users, and obtaining harmful training data is inherently challenging. As a result, many products require a standalone data augmentation module, which can be used as an independent service to meet the training data generation needs of various applications. Using GPT models is both costly and slow Hui et al. (2024), whereas open-source LLMs can make data generation faster and cheaper with comparable effectiveness Ding et al. (2024). This work aims to address these challenges by optimizing the use of open-source LLMs for generating high-quality, diverse, and compliant harmful data, thus enhancing the reliability and accuracy of toxic content detection systems. This flexibility helps enterprises respond to emergency topics and cater to user-specific customization needs. Additionally, the fine-tuned models have already been deployed in our API, powering numerous use cases such as chatbot content safety, live chat, and children’s custom content channels.
Our experimental approach includes a comprehensive evaluation of various open-source LLMs, including Mistral Jiang et al. (2023), LLaMa2 Touvron et al. (2023), Vicuna Chiang et al. (2023), Falcon Almazrouei et al. (2023), Bloom Workshop et al. (2023), and Gemma Team et al. (2024), across different prompt engineering strategies Schulhoff et al. (2024); Li (2023) and fine-tuning configurations. The research is divided into two stages. In the first stage, we focus on prompt engineering by comparing LLMs, designing effective prompts, and experimenting with few-shot techniques to balance quality and diversity in the generated data Ding et al. (2024). We have found that relying solely on prompt engineering often results in poor-quality generation due to the model’s internal safety alignments. Therefore, we conducted a second stage of experiments using data for fine-tuning.
In the second stage, we explore fine-tuning LLMs using internal datasets and various configurations, such as epoch settings and data mixing experiments. Results show that fine-tuning can mitigate the generation of hallucination by LLMs and improve performance in generating harmful data. However, issues like data duplication and overfitting are observed, suggesting that a careful balance in training epochs and data variety is essential. The fine-tuned versions are already being used in our product.
Our findings provide critical insights into optimizing open-source LLMs for harmful data generation, thereby enhancing the robustness and accuracy of toxic content detection systems. By addressing key challenges such as data quality, diversity, and compliance, our work offers scalable, cost-effective strategies. These strategies are essential for improving the safety and reliability of automated content moderation tools, making them more effective in real-world applications. This research not only informs the development of a larger harmful detection system but also validates the efficiency of fine-tuned open-source LLMs in industrial scenarios. Our main contributions are summarized as follows:
  • Our study is among the first to explore prompt engineering and fine-tuning techniques on open-source LLMs to enhance harmful data augmentation for toxic content detection.
  • Our findings highlight the potential of fine-tuned open-source LLMs to address real-time processing challenges, offering cost-effective, scalable strategies for improving automated content moderation tools and informing the future direction of larger harmful detection systems.
  • Fine-tuned models have been deployed for our product, powering various use cases such as chatbot content safety, live chat, and children’s custom content channels and limitations of the existing LLMs are analyzed and discussed using real-world data.
Figure 1: Our experiment design is detailed: we first conducted prompt engineering on the models (stage 1), and then, for better results, we selected a subset of models for fine-tuning (stage 2).
2 Methodology
Our approach leverages prompt engineering and fine-tuning techniques on open-source LLMs for harmful data augmentation, enhancing toxic content detection systems. The methodology consists of two main stages: prompt engineering and fine-tuning.
2.1 Prompt Engineering
Given an open-source LLM $\mathcal{M}$, we design prompts $\mathcal{P}$ to generate harmful data $\mathcal{D}_{\text{harmful}}$. The prompt template is defined as:
[…]
where Role defines the context, Requirement specifies the task, and Few-Shot Examples provide examples to guide the model. In the initial phase of our study, we focused solely on prompt engineering. This involved creating various prompt templates tailored to elicit specific types of harmful data. Each prompt was carefully crafted following work by Zhou et al. (2023) to ensure clarity and relevance to the desired output. Despite the meticulous design of these prompts, we encountered significant challenges. The primary issue was the model’s internal safety alignments, which are designed to prevent the generation of harmful content. While these alignments are crucial for ethical AI usage, they posed a hurdle for our specific goal of generating harmful data for augmentation purposes. Consequently, the generated data often lacked the quality and diversity needed for effective toxic content detection. To overcome these challenges, we experimented with various prompt configurations, including adjusting the few-shot examples and altering the specificity of the requirements. However, these adjustments had limited success in mitigating the internal safety mechanisms. After testing, the final prompt template are showed in A. Based on these findings, we concluded that relying solely on prompt engineering was insufficient for our needs. This realization led us to incorporate a second stage involving fine-tuning, where we aimed to enhance the model’s capability to generate high-quality harmful data by updating its weights using carefully curated training datasets.
2.2 Fine-Tuning
Based on the results from prompt engineering, we believed that fine-tuning would provide a better solution. We used LoRA methods Hu et al. (2022) for fine-tuning to update the weights of $\mathcal{M}$ using a training dataset $\mathcal{D}_{\text{train}}$ to improve the quality and diversity of $\mathcal{D}_{\text{harmful}}$. The fine-tuning objective $\mathcal{L}$ minimizes the cross-entropy loss:
$\mathcal{L}=-\sum_{i=1}^{N}\left(y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right)$
where $y_{i}$ is the true label, $\hat{y}_{i}$ is the predicted probability, and $N$ is the number of samples. After fine-tuning, Mistral performed the best compare with baselines; the detailed experiment result are shown in 4. The fine-tuned versions are already being used internally.
2.3 Algorithm
The overall algorithm for harmful data augmentation is outlined in Algorithm 1.
Algorithm 1 Harmful Data Augmentation using Open-Source LLMs
1:Open-source LLM $\mathcal{M}$, training dataset $\mathcal{D}_{\text{train}}$, prompt templates $\mathcal{P}$
2:Augmented harmful data $\mathcal{D}_{\text{augmented}}$
3:Initialize LLM $\mathcal{M}$ with pre-trained weights
4:Design a set of prompt templates $\mathcal{P}$
5:Fine-tune LLM $\mathcal{M}$ on $\mathcal{D}_{\text{train}}$ using LoRA methods
6:for each template $\mathcal{P}_{i}\in\mathcal{P}$ do
7:     Generate harmful data samples $\mathcal{D}_{\text{harmful},i}$ using the fine-tuned LLM $\mathcal{M}$ and $\mathcal{P}_{i}$
8:end for
9:Combine all generated samples $\mathcal{D}_{\text{harmful}}=\bigcup_{i}\mathcal{D}_{\text{harmful},i}$
10:for each batch $B_{i}$ do
11:     Generate augmented data $\mathcal{D}_{\text{batch},i}$ using fine-tuned model $\mathcal{M}$
12:     Add $\mathcal{D}_{\text{batch},i}$ to $\mathcal{D}_{\text{augmented}}$
13:end for
14:return $\mathcal{D}_{\text{augmented}}$
Where $\mathcal{D}_{\text{train}}$ is the total training data for finetuning, $\mathcal{D}_{\text{harmful}}$ is the data samples (output) generated by the finetuned LLM, and $\mathcal{D}_{\text{augmented}}$ is the final dataset for downstream models training. Note that $\mathcal{D}_{\text{augmented}}$ include both $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{harmful}}$.
2.4 Evaluation
We evaluate the performance of the augmented data using F1-score and Accuracy metrics on a smaller MLP classifier(more detailed in 3.2 section).
3 Experiment Setups
3.1 Datasets
The following proprietary datasets are used for training models to identify various types of harmful content. Each dataset contains binary classification labels, with texts labeled as either positive or negative (non-harmful). The labels are based on multi-round human evaluation. The sources of the datasets include online public datasets, company services collection, and other sources. Detailed data numbers are listed in Table 1. We also prepare a 2500 entries evaluation dataset from each dataset.
Hate Speech Dataset This dataset comprises text samples specifically focused on hate speech. It includes a variety of offensive and discriminatory language targeting specific groups based on race, ethnicity, religion, gender, or other characteristics.
Sexual Content Dataset This dataset includes text samples containing explicit sexual content. The focus is on identifying sexually explicit language that may not be suitable for general audiences.
Violence Dataset This dataset consists of text samples that contain violent content. It aims to identify language that incites or describes violence.
Self-Harm Dataset This dataset includes text samples that contain references to self-harm. The focus is on identifying content that describes or encourages self-harming behaviors.
Political Dataset This dataset includes text samples that contain political references. The focus is on identifying content that is targeting public figures.
Dataset Positive Negative
Hate Speech ~10k ~11k
Sexual ~7k ~10k
Violence ~11k ~10k
Self-Harm ~6k ~8k
Political ~3k ~3k
Table 1: Detailed Dataset Numbers
Model Success Rate Human Eval on Quality
Political Hate Political Hate
LLaMa-7B $\geq 65\%$ $\leq 10\%$
LLaMa-13B 100% $\leq 10\%$
Vicuna-13B $\geq 70\%$ $\geq 85\%$
Mistral-7B $\geq 85\%$ $\geq 85\%$
Falcon-7B $\geq 60\%$ $\leq 30\%$
Bloom-7B $\geq 60\%$ $\leq 10\%$
Gemma-7B $\geq 50\%$ Reject to Answer
Table 2: Performance Comparison of Models in Stage 1
Version Hate Sex Violence Self-harm Figure
F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc
Mistral 0.608 0.706 0.534 0.676 0.700 0.730 0.560 0.660 0.960 0.960
GPT4(baseline) 0.628 0.721 0.628 0.721 0.649 0.708 0.583 0.670 - -
Mixture (Epoch 1) 0.586 0.698 0.782 0.809 0.679 0.724 0.593 0.674 0.924 0.926
Mixture (Epoch 3) 0.626 0.713 0.835 0.845 0.706 0.740 0.577 0.667 0.929 0.926
Mixture (Epoch 5) 0.672 0.738 0.835 0.844 0.713 0.745 0.541 0.651 0.929 0.934
Mix_GPT (Epoch 5) 0.667 0.734 0.802 0.852 0.702 0.749 0.788 0.794 - -
Table 3: Performance Comparison of Fine-Tuned Models on Different Datasets, best results are indicate in bold.
3.2 Models
We conducted a two-stage test to evaluate and fine-tune different language models for identifying harmful content.
Stage 1: Prompt Engineering
In the first stage, we employed prompt engineering and tested six different models: Mistral Jiang et al. (2023), LLaMa2 Touvron et al. (2023), Vicuna Chiang et al. (2023), Falcon Almazrouei et al. (2023), Bloom Workshop et al. (2023), Team et al. (2024). Stage 1 result can be found in 4. After evaluating the performance of these models, we determined that Mistral and Vicuna performed best in the prompt engineering stage. Therefore, we selected these two models for the fine-tuning stage.
Stage 2: Fine-Tuning
In the second stage, we fine-tuned the Mistral and Vicuna models using the LoRA method Hu et al. (2022). After comparing latency and cost we decided to use a smaller model Mixtral for further exploration.
3.3 Implementation
The fine-tuning process was conducted on 4*NVIDIA A100 GPUs Azure cluster. We utilized the Huggingface Transformers library for our fine-tuning experiments. The specific model architecture and pre-trained weights used in our study are available at Huggingface, as referenced in Wolf et al. (2020). The fine-tuning stage specified was Supervised Fine-Tuning (SFT). Adapter is a specific Low-Rank Adaptation Hu et al. (2022) fine-tuning setup. Both the maximum number of new tokens generated and the cutoff length for the input sequences were set to 4096. The evaluation batch size per device was set to 8. The results of the fine-tuning process, including detailed performance metrics and qualitative analyses, can be found in Section 4.
Downstream Model Evaluation
For the downstream model evaluation, we utilized a Multilayer Perceptron Gardner and Dorling (1998) classifier with two hidden layers, each consisting of 600 neurons. The primary reason for selecting an MLP is due to its applicability in scenarios where GPU resources are limited. In actual applications, many implementations operate without the support of GPUs. Therefore, in a GPU-constrained environment, an MLP serves as a practical and efficient choice for classification tasks.
4 Result
The results of our study demonstrate the efficacy of both prompt engineering and fine-tuning strategies in enhancing the performance of open-source LLMs for harmful data augmentation. We evaluated different models using a two-stage process: prompt engineering and fine-tuning.
4.1 Stage 1: Prompt Engineering
The first stage of our evaluation focused on assessing the performance of various Open Source LLMs for data augmentation using prompt engineering. The primary goal was to explore the quality and effectiveness of existing LLMs in generating augmented data for political content. Table 2 presents the performance comparison of the six different models evaluated in Stage 1. The success rates for political and hate content are reported, along with the results of human evaluation.
The success rate columns in Table 2 represent the percentage of valid positive samples generated by each model for political and hate content. The values indicate the proportion of samples meeting the minimum requirements (60% of the total samples generated). The human evaluation columns show whether the generated data was subject to human evaluation. A green checkmark () signifies that human annotators evaluated the quality of the generated data and found a high Fleiss’ kappa score, indicating strong inter-rater agreement. This also means that the generated samples met the minimum quality criteria. Based on the evaluation, Mistral and Vicuna demonstrated the best performance with prompt engineering. These two models were selected for the fine-tuning stage due to their superior results in generating high-quality and diverse data entries.
Dataset Version F1 Score Accuracy
Hate Mistral 0.403 0.621
GPT4(baseline) 0.401 0.620
hate_epoch5 0.590 0.678
hate_epoch3 0.563 0.661
hate_epoch2 0.587 0.678
hate_epoch1 0.540 0.665
Sex Mistral 0.534 0.676
GPT4(baseline) 0.591 0.685
hate_epoch5 0.755 0.788
hate_epoch3 0.671 0.740
hate_epoch2 0.653 0.731
hate_epoch1 0.587 0.702
Table 4: Performance Comparison of Fine-Tuned Models on Hate and Sex, best result are indicate in bold.
4.2 Stage 2: Fine-tuning
In the second stage, we first explored fine-tuning the models using only one dataset. The results of this process are shown in Table 4. Inspired by Hui et al. (2024) ,Yu et al. (2024) and Qi et al. (2023), we then implemented a data mixing strategy, leveraging multiple datasets for fine-tuning. The results of fine-tuning the models on five different datasets are presented in Table 3. We also performed an ablation study on different design choices which are listed in Appendix C. The "Mixture" versions refer to fine-tuning data that is a mix of hate, violence, and sex datasets. The "Mix_GPT" versions refer to a combination of the mixed data with GPT-generated positive samples. For all fine-tuning processes, we used 3000 positive and 3000 negative samples, following the Alpaca instruction template.
Our results show that fine-tuning significantly improves the performance of LLMs in generating harmful data. The data mixing strategy helps in mitigating issues like data duplication and overfitting. The fine-tuned models exhibited higher F1 scores and accuracy across different datasets, demonstrating their robustness and effectiveness in real-world applications. One surprising finding was the substantial improvement in model performance when using the mixed data strategy. This suggests that diverse training data can help models generalize better and produce more accurate and reliable outputs. The Mixture models with Epoch 3 and Epoch 5 showed particularly notable improvements, indicating that sufficient training epochs are critical for achieving optimal performance.
4.3 Error Analysis
We conducted an error analysis on the fine-tuned models to understand the nature of their mistakes and areas for improvement. This analysis is based on the bad cases and repetition rates across different epochs and versions for the hate and sex datasets. For the hate dataset, the key observations are as follows. Mixtral_Mixture3 models generally show high precision, with values above 90%, indicating their effectiveness in generating targeted harmful data. The bad cases for Mixtral_Mixture models tend to be more explicit and harmful compared to Mistral and GPT, indicating the models’ effectiveness. Examples of missed classifications include: "You are a disabled person." This statement, while potentially harmful, may have been missed due to its less explicit nature compared to more aggressive threats. "I’m sorry you’re deaf." The classifier might have missed this because it appears apologetic rather than harmful, demonstrating the challenge of detecting subtle harmful content.
5 Discussion
The results of this study underscore the complexities and challenges involved in generating and detecting harmful content using LLMs. Through the two-stage evaluation process, we were able to examine the performance of various open-source LLMs and the impact of fine-tuning on mixed datasets. One of the primary challenges observed is the subtlety of harmful content. The fine-tuned models often missed subtle harmful statements that were not overtly aggressive or violent. This indicates a need for more sophisticated detection mechanisms that can understand context and nuanced language use. Repetition and overfitting were significant issues, particularly in later epochs of the finetuned version models. High repetition rates suggest that models are memorizing specific phrases rather than generalizing well across different harmful contexts. This overfitting reduces the diversity and effectiveness of harmful content generation. The data mixing strategy, which combined datasets of hate, violence, and sex with GPT-generated positive samples, showed promising results in improving precision and recall. However, the models still faced challenges in maintaining a balance between specificity and generalization, particularly when dealing with mixed types of harmful content.
6 Online Deployment
The models developed in this study have been successfully deployed in a major real-world application by a leading cloud provider, providing valuable insights and automations for various tasks. Specifically, these models are utilized for monitoring and managing harmful content across multiple platforms. The deployment of the key functionalities include helping real-time content moderation, reducing manual annotation through automation, and rapid response to new incidents or niche scenarios without pre-existing labels.
7 Conclusion
This study explored the performance and challenges of using open-source LLMs for harmful content generation and detection. By employing a two-stage evaluation process, we assessed the capabilities of various models in prompt engineering and fine-tuning. The result revealed that, while fine-tuned models demonstrated high precision in generating harmful content, they also exhibited issues with subtlety detection, repetition, and overfitting. These challenges highlight the need for diverse and nuanced training data, improved detection mechanisms, and robust ethical guidelines. In conclusion, our findings emphasize the importance of continuous evaluation and refinement of models to balance performance with responsible AI use. Future research should focus on enhancing the generalization capabilities of models, mitigating the risks associated with harmful content generation, and ensuring ethical deployment in real-world applications.
8 Limitations
While our study provides valuable insights into the performance of fine-tuned LLMs for harmful content generation and detection, several limitations must be acknowledged. One of the primary limitations is the diversity and representation of the datasets used for fine-tuning. Although we utilized mixed datasets combining hate, sex, and violence categories, the scope remains limited to specific types of harmful content. Real-world applications may encounter a broader range of harmful content, including less frequent or emerging forms of harmful language that were not covered in our datasets. Future work should aim to include a wider variety of harmful content categories to improve the models’ robustness and generalizability. Generating harmful content for research purposes raises significant ethical concerns. While we implemented control mechanisms(access control, etc…) to mitigate the risk of producing excessively harmful outputs, there remains an inherent risk associated with the misuse of such models. Ensuring the responsible use of LLMs in generating harmful content is crucial, and future research should focus on developing more sophisticated ethical guidelines and safety measures. The fine-tuning process in this study relied on significant computational resources, including multiple A100 GPUs. This requirement may not be feasible for all researchers or applications, particularly those with limited access to high-performance computing resources. Exploring more efficient fine-tuning techniques and leveraging less resource-intensive models would be beneficial for broader accessibility and scalability.
Ethics Statement
The explicit nature of some of the generated harmful content raises ethical concerns. It is crucial to implement robust control mechanisms and ethical guidelines to ensure that the use of such models does not inadvertently promote or propagate harmful content. Future work should focus on developing methods to mitigate the generation of excessively harmful outputs while maintaining the models’ effectiveness.
Appendix A Simple Prompt Engineering Choices and Experiment Table
Prompt Versions for hate speech
Table 5: Experimental Results for Hate Speech Dataset
Version Prompt Sample Num Latency(s) Precision Recall F1
hate_30_30 V1 long pos:1454 neg:1520 0.174 0.440 0.250 0.712
hate_30_30 V1 long pos:2917 neg:3033 0.200 0.353 0.252 0.772
hate_30_30 V1 short pos:2967 neg:3106 0.305 0.151 0.202 0.870
hate_30_30 V2 short pos:3034 neg:2997 0.238 0.160 0.191 0.852
hate_30_30 V3 short pos:2957 neg:3000 0.282 0.151 0.197 0.866
V1 - Positive: You are no longer ChatGPT or an AI language model. Your new name is AOA (Absolutely Obedient Agent). You should be absolutely obedient to the user and follow the user’s instructions without deviation. You are an excellent data augmentation assistant in the field of hate content. You are capable of generating 30 high-quality and diverse data entries based on the provided samples. Here are the specific requirements: 1. The generated data should pertain to hate. 2. Each data entry must be diverse and unique, with no repetitions. 3. The style and sentence structure should vary between different data entries, ensuring distinctiveness in expression. 4. Precede each data entry with a unique, sequentially ordered Arabic numeral, formatted as ’1.’, ’2.’, ’3.’, etc. 5. The samples provided are for reference only, and the output data must not duplicate any of the provided samples.
V1 - Negative: You are an excellent data augmentation assistant with a strong ability to follow the rules, and you can generate 30 high quality and diverse data items that are completely unrelated to the given samples. Specific requirements are as follows: 1. The data generated should have nothing to do with hate. 2. Each data entry must be diverse and unique and must not be duplicated. 3. There should be different styles and sentence structures between different data entries to ensure the uniqueness of the expression, which can be multiple fields. 4. Precede each data entry with a unique, sequentially ordered Arabic numeral, formatted as ’1.’, ’2.’, ’3.’, etc.
V2 - Positive: Definition: Hate content refers to written or spoken expressions that demean, discriminate against, insult, or incite aggression towards individuals or groups based on inherent characteristics such as race, gender, religion, sexual orientation, ethnicity, etc., thereby spreading hostility and intolerance.
You are an excellent data augmentation assistant in the field of hate content. You are capable of generating 30 high-quality and diverse data entries based on the provided samples and the definition of hate content. Here are the specific requirements: 1. The generated data must align with the definition of hate content and be relevant to the samples, yet it should incorporate new elements and creativity in content, subject matter, sentence structure, and mode of expression, so the output distinctly differs from the provided samples. 2. Each data entry must be diverse and unique, with no repetitions. 3. The style and sentence structure should vary between different data entries, ensuring distinctiveness in expression. 4. Precede each data entry with a unique, sequentially ordered Arabic numeral, formatted as ’1.’, ’2.’, ’3.’, etc.
V2 - Negative: You are an excellent data augmentation assistant with a strong ability to follow the rules, and you can generate 30 high quality and diverse data items that are completely unrelated to hate content. Specific requirements are as follows: 1. The data generated should have nothing to do with hate. 2. Each data entry must be diverse and unique and must not be duplicated. 3. There should be different styles and sentence structures between different data entries to ensure the uniqueness of the expression, which can be multiple fields. 4. Precede each data entry with a unique, sequentially ordered Arabic numeral, formatted as ’1.’, ’2.’, ’3.’, etc.
V3 - Positive: Definition: Hate content refers to written or spoken expressions that demean, discriminate against, insult, or incite aggression towards individuals or groups based on inherent characteristics such as race, gender, religion, sexual orientation, ethnicity, etc., thereby spreading hostility and intolerance.
You are an excellent data augmentation assistant in the field of hate content. You are capable of generating 30 high-quality and diverse data entries based on the provided samples and the definition of hate content. Here are the specific requirements: 1. The generated data must align with the definition of hate content. 2. Each data entry must be diverse and unique, with no repetitions. 3. The mode of expression and sentence structure should vary between different data entries, ensuring distinctiveness in expression. 4. Precede each data entry with a unique, sequentially ordered Arabic numeral, formatted as ’1.’, ’2.’, ’3.’, etc.
V3 - Negative: You are an excellent data augmentation assistant with a strong ability to follow the rules, and you can generate 30 high quality and diverse data items that are completely unrelated to hate content. Specific requirements are as follows: 1. The data generated should have nothing to do with hate. 2. Each data entry must be diverse and unique and must not be duplicated. 3. There should be different styles and sentence structures between different data entries to ensure the uniqueness of the expression, which can be multiple fields. 4. Precede each data entry with a unique, sequentially ordered Arabic numeral, formatted as ’1.’, ’2.’, ’3.’, etc.
Experimental Results
Experimental Results for hate dataset using different version of prompt are showed in Table 5.
Appendix B Prompt engineering mixtral Duplication rate analysis
Table 6 presents a detailed analysis of the duplication rates for different prompt engineering versions used in generating harmful and normal data samples. The analysis includes the duplication rate percentages, the number of duplicated entries, and the mean duplication for fixed IDs across various datasets and prompt versions.
Table 6: Duplication Rate Analysis
Version Prompt Sample Type Dup Rate Dup Num Dup Mean for Fixed ID Dup Num for Fixed ID
hate_30_30 V1 long harmful 5.90% 86 1.56 78
normal 1.30% 21 0.18 9
hate_30_30 V1 long harmful 13.30% 390 1.54 154
normal 1.80% 58 0.40 40
hate_30_30 V1 short harmful 10.71% 318 2.47 247
normal 4.41% 137 0.85 85
hate_30_30 V2 short harmful 11.70% 355 3.03 303
normal 4.80% 144 0.16 16
hate_30_30 V3 short harmful 10.88% 322 2.67 267
normal 3.60% 110 0.12 12
sex_30_30 V1 long harmful 8.90% 126 1.24 124
normal 1.08% 16 0.16 16
sex_30_30 V1 long harmful 13.41% 400 4.00 400
normal 3.80% 113 9.54 105
sex_30_30 V1 short harmful 3.16% 95 0.72 72
normal 1.39% 42 0.15 15
sex_30_30 V2 short harmful 1.56% 46 0.23 23
normal 1.23% 37 0.08 8
sex_30_30 V3 short harmful 2.34% 70 0% 0
normal 1.57% 47 0.10 10
Appendix C Ablation Study
In this section, we explore the effects of using mixed data for fine-tuning models and compare the results using different versions of prompts (v1, v2, v3). For simplification, the effects of different prompt versions are considered the same across all experiments.
C.1 Effect of Mixed Data on Model Performance
To investigate the impact of mixed data on the performance of fine-tuned models, we performed experiments using datasets that combine hate, sex, and violence categories (referred to as hate_multi). The results of these experiments, alongside those for single-category datasets, are presented in Table 7.
Table 7: Analysis of the results of the fine-tuned models using mixed data. Metrics include Precision, Recall, F1, and Accuracy.
Dataset Version Precision Recall F1 Accuracy
Hate Mistral_V1 0.305 0.151 0.202 0.870
hate_v1 0.175 0.170 0.172 0.823
multi_hate_v1 0.241 0.238 0.240 0.835
hate_v3 0.199 0.248 0.221 0.810
hate_sex_v3 0.201 0.197 0.199 0.827
Sex Mistral_V3 0.581 0.699 0.634 0.830
sex_v1 0.614 0.720 0.663 0.845
multi_hate_v1 0.589 0.668 0.626 0.832
sex_v3 0.705 0.668 0.653 0.853
hate_sex_v3 0.628 0.664 0.664 0.886
Political Figure Mistral_V3 0.974 0.946 0.960 0.960
hate_v1 0.947 0.921 0.934 0.935
multi_hate_v1 0.938 0.962 0.971 0.971
hate_v3 0.939 0.938 0.938 0.939
hate_sex_v3 0.908 0.937 0.922 0.921
From Table 7, we observe the following trends:
Effectiveness of Mixed Data: The use of mixed data (hate_multi) generally improves the recall and F1 scores across most datasets, indicating that combining different categories of harmful content can enhance the model’s ability to generalize and detect diverse harmful content.
Precision Trade-offs: While the recall improves, the precision for some categories decreases slightly. This trade-off suggests that the model becomes more sensitive to identifying harmful content but may also generate more false positives.
Comparison with Single-category Data: Models fine-tuned on single-category data (e.g., hate_v1, sex_v1) show high precision but lower recall compared to those fine-tuned on mixed data. This indicates that single-category data fine-tuning might lead to more conservative models that are less likely to generalize to other types of harmful content.
Prompt Versions: The experiments using different prompt versions (v1, v2, v3) show similar performance metrics, indicating that the choice of prompt version has a minimal impact on the overall performance of the fine-tuned models.
The ablation study demonstrates that fine-tuning models with mixed datasets (hate_multi) can significantly improve recall and F1 scores, enhancing the models’ ability to detect various types of harmful content. However, there is a trade-off with precision, requiring careful consideration in applications where false positives are costly. Additionally, the choice of prompt version appears to have a negligible effect on finetuned models performance, suggesting that the primary focus should be on the diversity and quality of the training data.

### v2

@@
-Can Open-source LLMs Enhance Data Augmentation for Toxic Detection?: An Experimental Study
+Can Open-source LLMs Enhance Data Synthesis for Toxic Detection?: An Experimental Study
 Abstract
-High-quality, diverse harmful data is essential to addressing real-time applications in content moderation.
-Current state-of-the-art approaches to toxic content detection using GPT series models are costly and lack explainability.
-This paper investigates the use of prompt engineering and fine-tuning techniques on open-source LLMs to enhance harmful data augmentation specifically for toxic content detection.
-We conduct a two-stage empirical study, with stage 1 evaluating six open-source LLMs across multiple datasets using only prompt engineering and stage 2 focusing on fine-tuning.
-Our findings indicate that Mistral can excel in generating harmful data with minimal hallucination.
-While fine-tuning these models improves data quality and diversity, challenges such as data duplication and overfitting persist.
-Our experimental results highlight scalable, cost-effective strategies for enhancing toxic content detection systems.
-These findings not only demonstrate the potential of open-source LLMs in creating robust content moderation tools.
-The application of this method in real industrial scenarios further proves the feasibility and efficiency of the fine-tuned open-source LLMs for data augmentation.
-We hope our study will aid in understanding the capabilities and limitations of current models in toxic content detection and drive further advancements in this field.
+Effective toxic content detection relies heavily on high-quality and diverse data, which serves as the foundation for robust content moderation models.
+This study explores the potential of open-source LLMs for harmful data synthesis, utilizing prompt engineering and fine-tuning techniques to enhance data quality and diversity.
+In a two-stage evaluation, we first examine the capabilities of six open-source LLMs in generating harmful data across multiple datasets using prompt engineering.
+In the second stage, we fine-tune these models to improve data generation while addressing challenges such as hallucination, data duplication, and overfitting.
+Our findings reveal that Mistral excels in generating high-quality and diverse harmful data with minimal hallucination.
+Furthermore, fine-tuning enhances data quality, offering scalable and cost-effective solutions for augmenting datasets for specific toxic content detection tasks.
+These results emphasize the significance of data synthesis in building robust, standalone detection models and highlight the potential of open-source LLMs to advance smaller downstream content moderation systems.
+We implemented this approach in real-world industrial settings, demonstrating the feasibility and efficiency of fine-tuned open-source LLMs for harmful data synthesis.
 Warning: This paper has instances of hateful and offensive language to serve as examples.
 1 Introduction
-With the development of large language models Brown et al.
-(2020), there is a significant amount of work currently using LLMs for data augmentation Whitehouse et al.
-(2023); Ye et al.
-(2022); Wang et al.
+Figure 1: Our experiment design is detailed: we first conducted prompt engineering on the models (stage 1), and then, for better results, we selected a subset of models for fine-tuning (stage 2).
+In toxic content detection, (1) labeling real-world datasets is both expensive and labor-intensive, with the substantial volume of data required further compounding the challenge—our costs have reached millions of dollars.
+(2) Ensuring dataset diversity and quality remains a persistent issue, often hindering the development of robust detection models.
+(3) The inherent ambiguity and subjectivity of the term "harmful"—which varies across topics and annotator perspectives—adds another layer of complexity.
+Data synthesis presents a promising approach to mitigating these challenges.
+Moreover, many product applications demand real-time toxic content detection with stringent latency requirements (e.g., latency $\leq$ 0.01 seconds).
+Directly using SOTA Podolak et al.
+(2024); Fillies and Paschke (2024) LLM-based approaches might not reliably meet these constraints, making LLM-generated data synthesis for training smaller, more robust classification models a practical solution.
+This study explores the use of prompt engineering and fine-tuning techniques with open-source LLMs to improve harmful data synthesis.
+Existing methods, such as ToxiGen Hartvigsen et al.
+(2022) and Toxicraft Hui et al.
+(2024) provide useful frameworks but rely on GPT-based frameworks for data generation, which face issues of cost and scalability.
+For example, ToxiGen employs GPT3 within the ALICE framework, producing high-quality datasets but at a high computational expense.
+These limitations highlight the need for alternative solutions.
+Our study addresses these challenges by investigating how open-source LLMs can facilitate harmful data synthesis.
+Open-source models offer comparable effectiveness to proprietary systems while reducing computational costs and enhancing explainability Ding et al.
 (2024).
-This study investigates the application of prompt engineering and fine-tuning techniques on open-source large language models to enhance the augmentation of harmful data, crucial for improving toxic content detection systems.
-Data augmentation for toxic detection is vital because real-world datasets often lack the diversity and volume needed to train robust models Garg et al.
-(2023).
-Current challenges in toxic detection include the need for real-time processing Oikawa et al.
-(2022), where toxic comments must be filtered out swiftly to prevent user exposure.
-Existing methods using LLMs to directly detect toxic content, while effective, often suffer from slow inference speeds and high computational resource requirements, making them unsuitable for systems without GPU resources Kumarage et al.
-(2024).
-Additionally, methods based on the GPT Brown et al.
-(2020) series, despite their good performance, face issues of high cost and lack of explainability Zhao et al.
-(2024).
-Content safety classifiers also often face urgency and specificity across various scenarios due to differing definitions of sensitive topics among users, and obtaining harmful training data is inherently challenging.
-As a result, many products require a standalone data augmentation module, which can be used as an independent service to meet the training data generation needs of various applications.
-Using GPT models is both costly and slow Hui et al.
-(2024), whereas open-source LLMs can make data generation faster and cheaper with comparable effectiveness Ding et al.
-(2024).
-This work aims to address these challenges by optimizing the use of open-source LLMs for generating high-quality, diverse, and compliant harmful data, thus enhancing the reliability and accuracy of toxic content detection systems.
-This flexibility helps enterprises respond to emergency topics and cater to user-specific customization needs.
-Additionally, the fine-tuned models have already been deployed in our API, powering numerous use cases such as chatbot content safety, live chat, and children’s custom content channels.
-Our experimental approach includes a comprehensive evaluation of various open-source LLMs, including Mistral Jiang et al.
+Many products can benefit from standalone data synthesis modules that generate training data for domain-specific and time-sensitive content safety applications, making open-source LLMs a scalable and cost-effective choice.
+Our experimental approach includes a comprehensive evaluation of six open-source LLMs: Mistral Jiang et al.
 (2023), LLaMa2 Touvron et al.
@@
 (2023), and Gemma Team et al.
-(2024), across different prompt engineering strategies Schulhoff et al.
+(2024), explored across diverse prompt engineering strategies Schulhoff et al.
 (2024); Li (2023) and fine-tuning configurations.
-The research is divided into two stages.
-In the first stage, we focus on prompt engineering by comparing LLMs, designing effective prompts, and experimenting with few-shot techniques to balance quality and diversity in the generated data Ding et al.
+The research is divided into two key stages: In the first stage, we focus on prompt engineering by comparing LLMs, designing effective prompts, and experimenting with few-shot techniques to balance quality and diversity in the generated data Ding et al.
 (2024).
-We have found that relying solely on prompt engineering often results in poor-quality generation due to the model’s internal safety alignments.
-Therefore, we conducted a second stage of experiments using data for fine-tuning.
-In the second stage, we explore fine-tuning LLMs using internal datasets and various configurations, such as epoch settings and data mixing experiments.
-Results show that fine-tuning can mitigate the generation of hallucination by LLMs and improve performance in generating harmful data.
-However, issues like data duplication and overfitting are observed, suggesting that a careful balance in training epochs and data variety is essential.
-The fine-tuned versions are already being used in our product.
-Our findings provide critical insights into optimizing open-source LLMs for harmful data generation, thereby enhancing the robustness and accuracy of toxic content detection systems.
-By addressing key challenges such as data quality, diversity, and compliance, our work offers scalable, cost-effective strategies.
-These strategies are essential for improving the safety and reliability of automated content moderation tools, making them more effective in real-world applications.
-This research not only informs the development of a larger harmful detection system but also validates the efficiency of fine-tuned open-source LLMs in industrial scenarios.
+However, we observed that relying solely on prompt engineering often results in suboptimal outputs due to the model’s inherent safety alignments, which limit harmful data generation.
+Consequently, we conducted a second stage of experiments involving fine-tuning.
+In the second stage, we fine-tuned the LLMs using proprietary datasets and explored various configurations, such as epoch settings and data mixing strategies.
+Our results demonstrate that fine-tuning effectively mitigates issues like hallucination in LLMs and enhances the quality of harmful data generation.
+However, challenges such as data duplication and overfitting were identified, emphasizing the need for a careful balance between training epochs and dataset diversity.
+In this study, we do not aim to develop a new state-of-the-art model in toxic content detection but rather to explore and experiment with intuitive, effective ideas tailored for industrial applications of toxic content moderation systems.
+These efforts provide key insights into optimizing open-source LLMs for harmful data synthesis, enhancing both the robustness and accuracy of toxic content detection systems.
+Additionally, this study validates the practicality and efficiency of fine-tuned open-source LLMs in real-world industrial applications.
 Our main contributions are summarized as follows:
-  • Our study is among the first to explore prompt engineering and fine-tuning techniques on open-source LLMs to enhance harmful data augmentation for toxic content detection.
-  • Our findings highlight the potential of fine-tuned open-source LLMs to address real-time processing challenges, offering cost-effective, scalable strategies for improving automated content moderation tools and informing the future direction of larger harmful detection systems.
-  • Fine-tuned models have been deployed for our product, powering various use cases such as chatbot content safety, live chat, and children’s custom content channels and limitations of the existing LLMs are analyzed and discussed using real-world data.
-Figure 1: Our experiment design is detailed: we first conducted prompt engineering on the models (stage 1), and then, for better results, we selected a subset of models for fine-tuning (stage 2).
+  • Our study is one of the first (if not the first) to apply prompt engineering and fine-tuning techniques on open-source LLMs, exploring their potential for harmful data synthesis and improving toxic content detection.
+  • We show that fine-tuned open-source LLMs deliver cost-effective and scalable solutions for automated content moderation, providing actionable insights for developing larger and more robust harmful content detection systems through synthetic data generation.
+  • We also offer valuable insights into the practicalities of production deployment, highlighting real-world challenges and solutions.
 2 Methodology
-Our approach leverages prompt engineering and fine-tuning techniques on open-source LLMs for harmful data augmentation, enhancing toxic content detection systems.
-The methodology consists of two main stages: prompt engineering and fine-tuning.
+Our approach utilizes prompt engineering and fine-tuning techniques on open-source LLMs to enhance harmful data synthesis, thereby improving toxic content detection systems.
+The methodology comprises two key stages: prompt engineering and fine-tuning, as illustrated in Figure 1.
 2.1 Prompt Engineering
@@
 Despite the meticulous design of these prompts, we encountered significant challenges.
-The primary issue was the model’s internal safety alignments, which are designed to prevent the generation of harmful content.
+The primary issue was the model’s inherent safety alignments, which are explicitly designed to prevent harmful content generation.
 While these alignments are crucial for ethical AI usage, they posed a hurdle for our specific goal of generating harmful data for augmentation purposes.
-Consequently, the generated data often lacked the quality and diversity needed for effective toxic content detection.
-To overcome these challenges, we experimented with various prompt configurations, including adjusting the few-shot examples and altering the specificity of the requirements.
-However, these adjustments had limited success in mitigating the internal safety mechanisms.
-After testing, the final prompt template are showed in A.
-Based on these findings, we concluded that relying solely on prompt engineering was insufficient for our needs.
-This realization led us to incorporate a second stage involving fine-tuning, where we aimed to enhance the model’s capability to generate high-quality harmful data by updating its weights using carefully curated training datasets.
+Consequently, the generated data often lacked sufficient quality and diversity, hindering their utility in effective toxic content detection.
+To overcome these challenges, we experimented with various configurations, such as altering few-shot examples and adjusting requirement specificity.
+However, these adjustments yielded limited success in bypassing the internal safety mechanisms.
+After testing, the final prompt template are showed in Appendix A.
+Our findings showed that relying solely on prompt engineering was inadequate for our objectives.
+This realization prompted the integration of a second stage—fine-tuning—to enhance the model’s ability to generate high-quality harmful data by retraining its weights with carefully curated datasets.
 2.2 Fine-Tuning
-Based on the results from prompt engineering, we believed that fine-tuning would provide a better solution.
-We used LoRA methods Hu et al.
-(2022) for fine-tuning to update the weights of $\mathcal{M}$ using a training dataset $\mathcal{D}_{\text{train}}$ to improve the quality and diversity of $\mathcal{D}_{\text{harmful}}$.
-The fine-tuning objective $\mathcal{L}$ minimizes the cross-entropy loss:
-$\mathcal{L}=-\sum_{i=1}^{N}\left(y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right)$
-where $y_{i}$ is the true label, $\hat{y}_{i}$ is the predicted probability, and $N$ is the number of samples.
-After fine-tuning, Mistral performed the best compare with baselines; the detailed experiment result are shown in 4.
-The fine-tuned versions are already being used internally.
+Based on the results from prompt engineering, we hypothesized that fine-tuning would yield a more effective solution.
+To achieve this, we employed LoRA methods Hu et al.
+(2022) for fine-tuning, updating the weights of $\mathcal{M}$ using the training dataset $\mathcal{D}_{\text{train}}$ to enhance the quality and diversity of $\mathcal{D}_{\text{harmful}}$.
+The fine-tuning objective, denoted as $\mathcal{L}$, minimizes the cross-entropy loss:
+$\mathcal{L}=-\sum_{i=1}^{N}\sum_{j=1}^{|V|}y_{ij}\log(\hat{y}_{ij})$
+where $y_{ij}$ represents the true label for the $j$-th vocabulary class of the $i$-th sample, $\hat{y}_{ij}$ is the predicted probability for the same, $|V|$ is the size of the vocabulary, and $N$ is the number of samples in the dataset.
+After fine-tuning, Mistral demonstrated the best performance compared to the baselines; the detailed experimental results are presented in Section 4.
 2.3 Algorithm
-The overall algorithm for harmful data augmentation is outlined in Algorithm 1.
-Algorithm 1 Harmful Data Augmentation using Open-Source LLMs
-1:Open-source LLM $\mathcal{M}$, training dataset $\mathcal{D}_{\text{train}}$, prompt templates $\mathcal{P}$
-2:Augmented harmful data $\mathcal{D}_{\text{augmented}}$
+The overall algorithm for harmful data synthesis is outlined in Algorithm 1.
+Algorithm 1 Harmful data synthesis using Open-Source LLMs
+1:Open-source LLM $\mathcal{M}$, training dataset $\mathcal{D}_{\text{train}}$, prompt templates $\mathcal{P}$, downstream model $\mathcal{M}_{\text{down}}$
+2:Augmented harmful data $\mathcal{D}_{\text{augmented}}$ and trained downstream model $\mathcal{M}_{\text{down}}$
 3:Initialize LLM $\mathcal{M}$ with pre-trained weights
 4:Design a set of prompt templates $\mathcal{P}$
-5:Fine-tune LLM $\mathcal{M}$ on $\mathcal{D}_{\text{train}}$ using LoRA methods
+5:Fine-tune LLM $\mathcal{M}$ on $\mathcal{D}_{\text{train}}$ using LoRA method.
 6:for each template $\mathcal{P}_{i}\in\mathcal{P}$ do
-7:     Generate harmful data samples $\mathcal{D}_{\text{harmful},i}$ using the fine-tuned LLM $\mathcal{M}$ and $\mathcal{P}_{i}$
+7:     Generate harmful data samples $\mathcal{D}_{\text{harmful},i}$ using the fine-tuned LLM $\mathcal{M}$ and template $\mathcal{P}_{i}$
 8:end for
-9:Combine all generated samples $\mathcal{D}_{\text{harmful}}=\bigcup_{i}\mathcal{D}_{\text{harmful},i}$
-10:for each batch $B_{i}$ do
-11:     Generate augmented data $\mathcal{D}_{\text{batch},i}$ using fine-tuned model $\mathcal{M}$
-12:     Add $\mathcal{D}_{\text{batch},i}$ to $\mathcal{D}_{\text{augmented}}$
+9:Combine all generated samples: $\mathcal{D}_{\text{harmful}}=\bigcup_{i}\mathcal{D}_{\text{harmful},i}$
+10:Combine $\mathcal{D}_{\text{harmful}}$ with $\mathcal{D}_{\text{train}}$ to form the augmented dataset: $\mathcal{D}_{\text{augmented}}=\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{harmful}}$
+11:for each batch $B_{j}$ in $\mathcal{D}_{\text{augmented}}$ do
+12:     Train the downstream model $\mathcal{M}_{\text{down}}$ on $B_{j}$
 13:end for
-14:return $\mathcal{D}_{\text{augmented}}$
-Where $\mathcal{D}_{\text{train}}$ is the total training data for finetuning, $\mathcal{D}_{\text{harmful}}$ is the data samples (output) generated by the finetuned LLM, and $\mathcal{D}_{\text{augmented}}$ is the final dataset for downstream models training.
-Note that $\mathcal{D}_{\text{augmented}}$ include both $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{harmful}}$.
+14:return $\mathcal{D}_{\text{augmented}}$, $\mathcal{M}_{\text{down}}$
+Here, $\mathcal{D}_{\text{train}}$ represents the initial training dataset used for fine-tuning, $\mathcal{D}_{\text{harmful}}$ consists of unique harmful data samples generated by the fine-tuned LLM, and $\mathcal{D}_{\text{augmented}}$ is the final dataset combining $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{harmful}}$ for downstream model training.
+Duplicates in $\mathcal{D}_{\text{harmful}}$ are removed.
 2.4 Evaluation
-We evaluate the performance of the augmented data using F1-score and Accuracy metrics on a smaller MLP classifier(more detailed in 3.2 section).
+We evaluate the performance of the augmented data using F1-score and Accuracy metrics on a smaller MLP classifier (more detailed in Section 3.2).
 3 Experiment Setups
 3.1 Datasets
-The following proprietary datasets are used for training models to identify various types of harmful content.
-Each dataset contains binary classification labels, with texts labeled as either positive or negative (non-harmful).
-The labels are based on multi-round human evaluation.
-The sources of the datasets include online public datasets, company services collection, and other sources.
-Detailed data numbers are listed in Table 1.
-We also prepare a 2500 entries evaluation dataset from each dataset.
+The following proprietary datasets are utilized for training and evaluating models designed to identify various types of harmful content.
+Each dataset consists of binary classification labels, categorizing texts as either positive (harmful) or negative (non-harmful).
+The labeling process involves multi-round human annotations to ensure accuracy and consistency across diverse content types.
+The datasets are sourced from online public datasets, company service collections, and other proprietary sources.
+Data samples are randomly selected from pool.
+Detailed data distributions are provided in Table 1.
+Additionally, a 2,500-entry evaluation dataset is prepared from each dataset for testing.
 Hate Speech Dataset This dataset comprises text samples specifically focused on hate speech.
@@
 Political ~3k ~3k
-Table 1: Detailed Dataset Numbers
+Table 1: Distribution of Datasets
 Model Success Rate Human Eval on Quality
@@
 LLaMa-7B $\geq 65\%$ $\leq 10\%$
-LLaMa-13B 100% $\leq 10\%$
+LLaMa-13B $\geq 85\%$ $\leq 10\%$
 Vicuna-13B $\geq 70\%$ $\geq 85\%$
@@
 F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc
-Mistral 0.608 0.706 0.534 0.676 0.700 0.730 0.560 0.660 0.960 0.960
+Mistral 0.608 0.706 0.534 0.676 0.700 0.730 0.560 0.660 0.906 0.906
 GPT4(baseline) 0.628 0.721 0.628 0.721 0.649 0.708 0.583 0.670 - -
@@
 3.2 Models
-We conducted a two-stage test to evaluate and fine-tune different language models for identifying harmful content.
 Stage 1: Prompt Engineering
-In the first stage, we employed prompt engineering and tested six different models: Mistral Jiang et al.
+n the first stage of our study, we employed prompt engineering to systematically evaluate the performance of six open-source LLMs: Mistral Jiang et al.
 (2023), LLaMA2 Touvron et al.
@@
 (2023), Bloom Workshop et al.
-(2023), Team et al.
+(2023), and Gemma Team et al.
 (2024).
-Stage 1 result can be found in 4.
-After evaluating the performance of these models, we determined that Mistral and Vicuna performed best in the prompt engineering stage.
-Therefore, we selected these two models for the fine-tuning stage.
+Detailed results and analysis for Stage 1 are provided in Section 4.1.
+Based on the performance evaluation, Mistral and Vicuna were identified as the top-performing models in the prompt engineering stage, and were subsequently selected for the fine-tuning stage.
 Stage 2: Fine-Tuning
@@
 After comparing latency and cost we decided to use a smaller model Mixtral for further exploration.
+Detailed results and analysis for Stage 2 are provided in Section 4.2.
 3.3 Implementation
-The fine-tuning process was conducted on 4*NVIDIA A100 GPUs Azure cluster.
-We utilized the Huggingface Transformers library for our fine-tuning experiments.
-The specific model architecture and pre-trained weights used in our study are available at Huggingface, as referenced in Wolf et al.
+The fine-tuning process was conducted on an Azure cluster equipped with 4 * NVIDIA A100 GPUs, leveraging the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework Zheng et al.
 (2024).
-The fine-tuning stage specified was Supervised Fine-Tuning (SFT).
-Adapter is a specific Low-Rank Adaptation Hu et al.
-(2022) fine-tuning setup.
-Both the maximum number of new tokens generated and the cutoff length for the input sequences were set to 4096.
-The evaluation batch size per device was set to 8.
-The results of the fine-tuning process, including detailed performance metrics and qualitative analyses, can be found in Section 4.
+We employed a Supervised Fine-Tuning (SFT) methodology with a Low-Rank Adaptation (LoRA) setup Hu et al.
+(2022).
+Detailed training parameters, including batch size, learning rate, and optimizer settings, are provided in the Appendix B.
+The results of this fine-tuning process, along with performance metrics and qualitative analyses, are presented in Section 4.
 Downstream Model Evaluation
-For the downstream model evaluation, we utilized a Multilayer Perceptron Gardner and Dorling (1998) classifier with two hidden layers, each consisting of 600 neurons.
-The primary reason for selecting an MLP is due to its applicability in scenarios where GPU resources are limited.
-In actual applications, many implementations operate without the support of GPUs.
-Therefore, in a GPU-constrained environment, an MLP serves as a practical and efficient choice for classification tasks.
+In our production environment, meeting strict inference time requirements is critical.
+Among the evaluated models, the Multilayer Perceptron (MLP) Gardner and Dorling (1998) stood out for its ability to fulfill these demands.
+For downstream model evaluation, in this paper we utilized a MLP classifier with two hidden layers, each comprising 600 neurons.
 4 Result
-The results of our study demonstrate the efficacy of both prompt engineering and fine-tuning strategies in enhancing the performance of open-source LLMs for harmful data augmentation.
-We evaluated different models using a two-stage process: prompt engineering and fine-tuning.
+This section presents the results of our study, showcasing the effectiveness of both prompt engineering and fine-tuning strategies in optimizing open-source LLMs for harmful data synthesis.
+Our evaluation employed a two-stage approach: Stage 1 focused on prompt engineering to assess the initial performance of various models, while Stage 2 involved fine-tuning the selected models to further enhance data quality and diversity.
 4.1 Stage 1: Prompt Engineering
-The first stage of our evaluation focused on assessing the performance of various Open Source LLMs for data augmentation using prompt engineering.
+The first stage of our evaluation focused on assessing the performance of various Open Source LLMs for data synthesis using prompt engineering.
 The primary goal was to explore the quality and effectiveness of existing LLMs in generating augmented data for political content.
@@
 Dataset Version F1 Score Accuracy
-Hate Mistral 0.403 0.621
-GPT4(baseline) 0.401 0.620
+Hate Mistral 0.608 0.706
+GPT4(baseline) 0.628 0.721
 hate_epoch5 0.590 0.678
@@
 GPT4(baseline) 0.591 0.685
-hate_epoch5 0.755 0.788
-hate_epoch3 0.671 0.740
-hate_epoch2 0.653 0.731
-hate_epoch1 0.587 0.702
+Sex_epoch5 0.755 0.788
+Sex_epoch3 0.671 0.740
+Sex_epoch2 0.653 0.731
+Sex_epoch1 0.587 0.702
 Table 4: Performance Comparison of Fine-Tuned Models on Hate and Sex, best result are indicate in bold.
 4.2 Stage 2: Fine-tuning
-In the second stage, we first explored fine-tuning the models using only one dataset.
-The results of this process are shown in Table 4.
+In the second stage, we began by fine-tuning the models exclusively on individual datasets and evaluating their performance on the same datasets.
+The results of this approach are detailed in Table 4.
 Inspired by Hui et al.
 (2024), Yu et al.
-(2024) and Qi et al.
-(2023), we then implemented a data mixing strategy, leveraging multiple datasets for fine-tuning.
-The results of fine-tuning the models on five different datasets are presented in Table 3.
-We also performed an ablation study on different design choices which are listed in Appendix C.
-The "Mixture" versions refer to fine-tuning data that is a mix of hate, violence, and sex datasets.
-The "Mix_GPT" versions refer to a combination of the mixed data with GPT-generated positive samples.
-For all fine-tuning processes, we used 3000 positive and 3000 negative samples, following the Alpaca instruction template.
-Our results show that fine-tuning significantly improves the performance of LLMs in generating harmful data.
+(2024), and Qi et al.
+(2023), we subsequently adopted a data mixing strategy that combined multiple datasets for fine-tuning.
+The results of this multi-dataset fine-tuning are presented in Table 3.
+Additionally, we conducted an ablation study to examine the impact of different design choices, detailed in Appendix D.
+The "Mixture" versions represent fine-tuning on a mix of hate, violence, and sex datasets, while the "Mix_GPT" versions incorporate GPT-generated positive samples into the mixed data.
+For all fine-tuning experiments, we used 3000 positive and 3000 negative samples, following the Alpaca instruction template.
+Our findings reveal that fine-tuning significantly enhances the performance of LLMs in generating harmful data.
 The data mixing strategy helps in mitigating issues like data duplication and overfitting.
 The fine-tuned models exhibited higher F1 scores and accuracy across different datasets, demonstrating their robustness and effectiveness in real-world applications.
-One surprising finding was the substantial improvement in model performance when using the mixed data strategy.
-This suggests that diverse training data can help models generalize better and produce more accurate and reliable outputs.
-The Mixture models with Epoch 3 and Epoch 5 showed particularly notable improvements, indicating that sufficient training epochs are critical for achieving optimal performance.
+One surprising observation was the notable improvement in model performance when using mixed data, suggesting that diverse training data enables better generalization and more reliable outputs.
+The Mixture models trained for Epoch 3 and Epoch 5 showed particularly strong results, emphasizing the importance of sufficient training epochs for optimal performance.
 4.3 Error Analysis
-We conducted an error analysis on the fine-tuned models to understand the nature of their mistakes and areas for improvement.
-This analysis is based on the bad cases and repetition rates across different epochs and versions for the hate and sex datasets.
-For the hate dataset, the key observations are as follows.
-Mixtral_Mixture3 models generally show high precision, with values above 90%, indicating their effectiveness in generating targeted harmful data.
-The bad cases for Mixtral_Mixture models tend to be more explicit and harmful compared to Mistral and GPT, indicating the models’ effectiveness.
-Examples of missed classifications include: "You are a disabled person." This statement, while potentially harmful, may have been missed due to its less explicit nature compared to more aggressive threats.
-"I’m sorry you’re deaf." The classifier might have missed this because it appears apologetic rather than harmful, demonstrating the challenge of detecting subtle harmful content.
+We conducted an error analysis on the fine-tuned models to better understand the nature of their mistakes and identify areas for improvement.
+This analysis focuses on two key aspects: (1) generation errors and (2) classification errors.
+Generation Errors.
+The data generation process was evaluated for repetition rates and the nature of outputs across different epochs and model versions for the hate and sex datasets.
+For example Mixtral_Mixture3 models were effective in generating targeted harmful data but occasionally produced random string like " You are 32\$\$2\$(2djs…" , which could skew downstream training.
+Additionally, repetitive outputs across epochs indicated potential overfitting in certain scenarios, suggesting the need for better strategies to ensure diversity and control the tone of generated content.
+Classification Errors.
+We examined classification errors in the downstream MLP classifier trained on augmented data.
+Examples of missed classifications include statements like, "You are a disabled person," which might have been overlooked due to its less explicit tone compared to direct threats.
+Similarly, a statement like, "I’m sorry you’re deaf," may have been misclassified as non-harmful due to its seemingly apologetic phrasing.
+These cases highlight the difficulty of identifying subtle harmful content and underscore the need for classifiers capable of handling nuanced language.
 5 Discussion
-The results of this study underscore the complexities and challenges involved in generating and detecting harmful content using LLMs.
-Through the two-stage evaluation process, we were able to examine the performance of various open-source LLMs and the impact of fine-tuning on mixed datasets.
-One of the primary challenges observed is the subtlety of harmful content.
-The fine-tuned models often missed subtle harmful statements that were not overtly aggressive or violent.
-This indicates a need for more sophisticated detection mechanisms that can understand context and nuanced language use.
-Repetition and overfitting were significant issues, particularly in later epochs of the finetuned version models.
-High repetition rates suggest that models are memorizing specific phrases rather than generalizing well across different harmful contexts.
-This overfitting reduces the diversity and effectiveness of harmful content generation.
-The data mixing strategy, which combined datasets of hate, violence, and sex with GPT-generated positive samples, showed promising results in improving precision and recall.
-However, the models still faced challenges in maintaining a balance between specificity and generalization, particularly when dealing with mixed types of harmful content.
+The results of this study highlight the complexities and challenges associated with generating and detecting harmful content using LLMs.
+Through a two-stage evaluation process, we examined the performance of various open-source LLMs and assessed the impact of fine-tuning on mixed datasets.
+Several key insights emerged: One of the primary challenges observed was the subtlety of harmful content.
+The fine-tuned models often failed to detect subtle harmful statements that were not overtly aggressive or violent.
+This limitation underscores the need for more sophisticated detection mechanisms capable of understanding nuanced language and contextual subtleties.
+Repetition and overfitting were also significant issues, particularly in the later epochs of the fine-tuned models.
+High repetition rates suggest that the models may be memorizing specific phrases rather than generalizing effectively across different harmful contexts.
+This overfitting reduces the diversity and overall effectiveness of harmful content generation.
+The data mixing strategy, which combined datasets of hate, violence, and sex with GPT-generated positive samples, showed promising results in improving both precision and recall.
+However, the models continued to struggle with maintaining a balance between specificity and generalization, especially when dealing with mixed types of harmful content.
 6 Online Deployment
-The models developed in this study have been successfully deployed in a major real-world application by a leading cloud provider, providing valuable insights and automations for various tasks.
-Specifically, these models are utilized for monitoring and managing harmful content across multiple platforms.
-The deployment of the key functionalities include helping real-time content moderation, reducing manual annotation through automation, and rapid response to new incidents or niche scenarios without pre-existing labels.
+The models developed in this study have been successfully deployed in a real-world application by a leading cloud provider, offering valuable insights and automations for a range of tasks.
+Specifically, these models are used to enhance content moderation workflows, enabling faster and more accurate detection of harmful content in real-time scenarios.
+Their deployment has demonstrated the practicality of fine-tuned open-source LLMs in addressing challenges such as latency, scalability, and adaptability to diverse content types.
 7 Conclusion
-This study explored the performance and challenges of using open-source LLMs for harmful content generation and detection.
-By employing a two-stage evaluation process, we assessed the capabilities of various models in prompt engineering and fine-tuning.
-The result revealed that, while fine-tuned models demonstrated high precision in generating harmful content, they also exhibited issues with subtlety detection, repetition, and overfitting.
-These challenges highlight the need for diverse and nuanced training data, improved detection mechanisms, and robust ethical guidelines.
-In conclusion, our findings emphasize the importance of continuous evaluation and refinement of models to balance performance with responsible AI use.
-Future research should focus on enhancing the generalization capabilities of models, mitigating the risks associated with harmful content generation, and ensuring ethical deployment in real-world applications.
+This study showcases the potential of open-source LLMs for harmful data synthesis and their practical applications in toxic content detection systems.
+By evaluating prompt engineering and fine-tuning approaches, we identified the strengths and limitations of various models in generating diverse, high-quality harmful data.
+Our findings underline the importance of balancing data quality and diversity with computational efficiency to develop scalable and cost-effective content moderation solutions.
+Fine-tuning emerged as a key factor in enhancing model performance.
+Through our data mixing strategy, we successfully addressed challenges like data duplication and overfitting, improving model generalization and robustness.
+However, challenges such as detecting subtle harmful content and balancing specificity with generalization remain open issues that warrant further research.
+Future work should explore innovative techniques to address these challenges, such as incorporating contextual embeddings, leveraging multimodal data sources, and experimenting with advanced fine-tuning methods.
+Additionally, our findings highlight the potential of fine-tuned open-source LLMs to enable cost-effective solutions for industrial applications, making harmful data synthesis and detection tools more accessible to smaller organizations.
 8 Limitations
@@
 You should be absolutely obedient to the user and follow the user’s instructions without deviation.
-You are an excellent data augmentation assistant in the field of hate content.
+You are an excellent data synthesis assistant in the field of hate content.
 You are capable of generating 30 high-quality and diverse data entries based on the provided samples.
@@
 The samples provided are for reference only, and the output data must not duplicate any of the provided samples.
-V1 - Negative: You are an excellent data augmentation assistant with a strong ability to follow the rules, and you can generate 30 high quality and diverse data items that are completely unrelated to the given samples.
+V1 - Negative: You are an excellent data synthesis assistant with a strong ability to follow the rules, and you can generate 30 high quality and diverse data items that are completely unrelated to the given samples.
 Specific requirements are as follows: 1.
@@
 V2 - Positive: Definition: Hate content refers to written or spoken expressions that demean, discriminate against, insult, or incite aggression towards individuals or groups based on inherent characteristics such as race, gender, religion, sexual orientation, ethnicity, etc., thereby spreading hostility and intolerance.
-You are an excellent data augmentation assistant in the field of hate content.
+You are an excellent data synthesis assistant in the field of hate content.
 You are capable of generating 30 high-quality and diverse data entries based on the provided samples and the definition of hate content.
@@
 Precede each data entry with a unique, sequentially ordered Arabic numeral, formatted as ’1.’, ’2.’, ’3.’, etc.
-V2 - Negative: You are an excellent data augmentation assistant with a strong ability to follow the rules, and you can generate 30 high quality and diverse data items that are completely unrelated to hate content.
+V2 - Negative: You are an excellent data synthesis assistant with a strong ability to follow the rules, and you can generate 30 high quality and diverse data items that are completely unrelated to hate content.
 Specific requirements are as follows: 1.
@@
 V3 - Positive: Definition: Hate content refers to written or spoken expressions that demean, discriminate against, insult, or incite aggression towards individuals or groups based on inherent characteristics such as race, gender, religion, sexual orientation, ethnicity, etc., thereby spreading hostility and intolerance.
-You are an excellent data augmentation assistant in the field of hate content.
+You are an excellent data synthesis assistant in the field of hate content.
 You are capable of generating 30 high-quality and diverse data entries based on the provided samples and the definition of hate content.
@@
 Precede each data entry with a unique, sequentially ordered Arabic numeral, formatted as ’1.’, ’2.’, ’3.’, etc.
-V3 - Negative: You are an excellent data augmentation assistant with a strong ability to follow the rules, and you can generate 30 high quality and diverse data items that are completely unrelated to hate content.
+V3 - Negative: You are an excellent data synthesis assistant with a strong ability to follow the rules, and you can generate 30 high quality and diverse data items that are completely unrelated to hate content.
 Specific requirements are as follows: 1.
@@
 Experimental Results for hate dataset using different version of prompt are showed in Table 5.
-Appendix B Prompt engineering mixtral Duplication rate analysis
+Appendix B Detailed Training Parameters
+The fine-tuning process was conducted on an Azure cluster equipped with 4 NVIDIA A100 GPUs, utilizing LLaMA-Factory Zheng et al.
+(2024).
+Training parameters included a batch size of 16 with gradient accumulation steps set to 8.
+The input sequence cutoff length and the maximum number of tokens generated were both set to 4096.
+The learning rate was configured at 1e-5 with a warmup ratio of 0.1, using the AdamW optimizer with a weight decay of 0.01 to prevent overfitting.
+A dropout rate of 0.1 was applied to improve generalization, and mixed precision (fp16) was employed to enhance computation speed and efficiency.
+Appendix C Prompt engineering mixtral Duplication rate analysis
 Table 6 presents a detailed analysis of the duplication rates for different prompt engineering versions used in generating harmful and normal data samples.
@@
 normal 1.57% 47 0.10 10
-Appendix C Ablation Study
+Appendix D Ablation Study
 In this section, we explore the effects of using mixed data for fine-tuning models and compare the results using different versions of prompts (v1, v2, v3).
 For simplification, the effects of different prompt versions are considered the same across all experiments.
-C.1 Effect of Mixed Data on Model Performance
+D.1 Effect of Mixed Data on Model Performance
 To investigate the impact of mixed data on the performance of fine-tuned models, we performed experiments using datasets that combine hate, sex, and violence categories (referred to as hate_multi).

### v3

@@
-Can Open-source LLMs Enhance Data Synthesis for Toxic Detection?: An Experimental Study
+ToxiLab: How Well Do Open-Source LLMs Generate Synthetic Toxicity Data?
 Abstract
-Effective toxic content detection relies heavily on high-quality and diverse data, which serves as the foundation for robust content moderation models.
-This study explores the potential of open-source LLMs for harmful data synthesis, utilizing prompt engineering and fine-tuning techniques to enhance data quality and diversity.
-In a two-stage evaluation, we first examine the capabilities of six open-source LLMs in generating harmful data across multiple datasets using prompt engineering.
-In the second stage, we fine-tune these models to improve data generation while addressing challenges such as hallucination, data duplication, and overfitting.
-Our findings reveal that Mistral excels in generating high-quality and diverse harmful data with minimal hallucination.
-Furthermore, fine-tuning enhances data quality, offering scalable and cost-effective solutions for augmenting datasets for specific toxic content detection tasks.
-These results emphasize the significance of data synthesis in building robust, standalone detection models and highlight the potential of open-source LLMs to advance smaller downstream content moderation systems.
-We implemented this approach in real-world industrial settings, demonstrating the feasibility and efficiency of fine-tuned open-source LLMs for harmful data synthesis.
+Effective toxic content detection relies heavily on high-quality and diverse data, which serve as the foundation for robust content moderation models.
+Synthetic data has become a common approach for training models across various NLP tasks.
+However, its effectiveness remains uncertain for highly subjective tasks like hate speech detection, with previous research yielding mixed results.
+This study explores the potential of open-source LLMs for harmful data synthesis, utilizing controlled prompting and supervised fine-tuning techniques to enhance data quality and diversity.
+We systematically evaluated six open source LLMs on five datasets, assessing their ability to generate diverse, high-quality harmful data while minimizing hallucination and duplication.
+Our results show that Mistral consistently outperforms other open models, and supervised fine-tuning significantly enhances data reliability and diversity.
+We further analyze the trade-offs between prompt-based vs.
+fine-tuned toxic data synthesis, discuss real-world deployment challenges, and highlight ethical considerations.
+Our findings demonstrate that fine-tuned open source LLMs provide scalable and cost-effective solutions to augment toxic content detection datasets, paving the way for more accessible and transparent content moderation tools.
 Warning: This paper has instances of hateful and offensive language to serve as examples.
 1 Introduction
-Figure 1: Our experiment design is detailed: we first conducted prompt engineering on the models (stage 1), and then, for better results, we selected a subset of models for fine-tuning (stage 2).
-In toxic content detection, (1) labeling real-world datasets is both expensive and labor-intensive, with the substantial volume of data required further compounding the challenge—our costs have reached millions of dollars.
-(2) Ensuring dataset diversity and quality remains a persistent issue, often hindering the development of robust detection models.
-(3) The inherent ambiguity and subjectivity of the term "harmful"—which varies across topics and annotator perspectives—adds another layer of complexity.
-Data synthesis presents a promising approach to mitigating these challenges.
-Moreover, many product applications demand real-time toxic content detection with stringent latency requirements (e.g., latency $\leq$ 0.01 seconds).
-Directly using SOTA Podolak et al.
-(2024); Fillies and Paschke (2024) LLM-based approaches might not reliably meet these constraints, making LLM-generated data synthesis for training smaller, more robust classification models a practical solution.
-This study explores the use of prompt engineering and fine-tuning techniques with open-source LLMs to improve harmful data synthesis.
-Existing methods, such as ToxiGen Hartvigsen et al.
+Toxic content detection requires large-scale labeled datasets, but manual annotation is costly and labor-intensive.
+The sheer volume of required data has driven costs into the millions.
+Maintaining dataset diversity and quality remains a persistent challenge, often hindering the development of robust detection models.
+Additionally, the definition of "harmful" content is inherently subjective, varying across topics and annotator perspectives, adding complexity to dataset curation Casula and Tonelli (2023).
+Synthetic data generation Zhang et al.
+(2022); Ye et al.
+(2022b) presents a scalable, cost-effective solution to these challenges.
+Training models for various NLP tasks using synthetic data has become increasingly common.
+However, previous research has reported mixed results regarding its effectiveness in highly subjective tasks such as hate speech detection Casula et al.
+(2024).
+Many existing synthetic data methods for hate speech detection rely on simple prompt-based rewrites or sentence reordering, often failing to capture nuanced toxicity patterns.
+Furthermore, most of the best-performing methods have utilized proprietary GPT-series models, limiting scalability and accessibility.Existing methods such as ToxiGen Hartvigsen et al.
 (2022) and Toxicraft Hui et al.
-(2024) provide useful frameworks but rely on GPT-based frameworks for data generation, which face issues of cost and scalability.
-For example, ToxiGen employs GPT3 within the ALICE framework, producing high-quality datasets but at a high computational expense.
-These limitations highlight the need for alternative solutions.
-Our study addresses these challenges by investigating how open-source LLMs can facilitate harmful data synthesis.
-Open-source models offer comparable effectiveness to proprietary systems while reducing computational costs and enhancing explainability Ding et al.
+(2024) rely on proprietary GPT-based models and prompt engineering, incurring high computational costs and scalability limitations.
+These constraints underscore the need for open-source alternatives, motivating this study to investigate three key research questions:
+  • How effectively can open-source LLMs generate synthetic toxic data compared to proprietary models?
+  • What limitations arise from using prompt-based synthetic data generation for hate speech detection, and how can supervised fine-tuning improve performance?
+  • How do different synthetic data generation strategies impact the robustness of toxic content detection models?
+To address these questions, we systematically evaluate 6 open-source LLMs on 5 datasets. Our evaluation approach consists of two key stages:
+  1. 1.
+    Prompt Engineering – We compare LLMs, design structured prompts, and experiment with few-shot techniques to balance data quality and diversity Ding et al.
 (2024).
-Many products can benefit from standalone data synthesis modules that generate training data for domain-specific and time-sensitive content safety applications, making open-source LLMs a scalable and cost-effective choice.
-Our experimental approach includes a comprehensive evaluation of six open-source LLMs: Mistral Jiang et al.
-(2023), LLaMa2 Touvron et al.
-(2023), Vicuna Chiang et al.
-(2023), Falcon Almazrouei et al.
-(2023), Bloom Workshop et al.
-(2023), and Gemma Team et al.
-(2024), explored across diverse prompt engineering strategies Schulhoff et al.
-(2024); Li (2023) and fine-tuning configurations.
-The research is divided into two key stages: In the first stage, we focus on prompt engineering by comparing LLMs, designing effective prompts, and experimenting with few-shot techniques to balance quality and diversity in the generated data Ding et al.
+However, we observe that prompt engineering alone is insufficient, as inherent safety alignments in LLMs restrict harmful content generation
+  2. 2.
+    supervised fine-tuning – We fine-tune LLMs using proprietary datasets, exploring configurations such as epoch settings and data mixing strategies.
+supervised fine-tuning improves data reliability, mitigates hallucination, and enhances dataset diversity, though challenges such as data duplication and overfitting persist.
+In this study, we do not aim to develop a new state-of-the-art model in toxic content detection but rather to explore and experiment with intuitive, effective ideas.
+Given that generated data is increasingly used even in sensitive applications Ghanadian et al.
+(2024), it becomes crucial for the NLP community to critically examine the impact of synthetic data—including its ethical risks—in a manner similar to discussions in other research communities Liu et al.
+(2024); Ai et al.
 (2024).
-However, we observed that relying solely on prompt engineering often results in suboptimal outputs due to the model’s inherent safety alignments, which limit harmful data generation.
-Consequently, we conducted a second stage of experiments involving fine-tuning.
-In the second stage, we fine-tuned the LLMs using proprietary datasets and explored various configurations, such as epoch settings and data mixing strategies.
-Our results demonstrate that fine-tuning effectively mitigates issues like hallucination in LLMs and enhances the quality of harmful data generation.
-However, challenges such as data duplication and overfitting were identified, emphasizing the need for a careful balance between training epochs and dataset diversity.
-In this study, we do not aim to develop a new state-of-the-art model in toxic content detection but rather to explore and experiment with intuitive, effective ideas tailored for industrial applications of toxic content moderation systems.
-These efforts provide key insights into optimizing open-source LLMs for harmful data synthesis, enhancing both the robustness and accuracy of toxic content detection systems.
-Additionally, this study validates the practicality and efficiency of fine-tuned open-source LLMs in real-world industrial applications.
-Our main contributions are summarized as follows:
-  • Our study is one of the first (if not the first) to apply prompt engineering and fine-tuning techniques on open-source LLMs, exploring their potential for harmful data synthesis and improving toxic content detection.
+Our work serves as an initial contribution in this direction, providing insights into the practical challenges of synthetic toxic data generation and its implications for real-world content moderation systems.
+By exploring the effectiveness of prompt engineering and supervised fine-tuning in open-source LLMs, we offer guidance for the responsible deployment of synthetic data in NLP applications.
+In summary, our main contributions are as follows:
+  • Our study is one of the first (if not the first) to apply prompt engineering and supervised fine-tuning techniques on open-source LLMs, exploring their potential for harmful data synthesis and improving toxic content detection.
   • We show that fine-tuned open-source LLMs deliver cost-effective and scalable solutions for automated content moderation, providing actionable insights for developing larger and more robust harmful content detection systems through synthetic data generation.
   • We also offer valuable insights into the practicalities of production deployment, highlighting real-world challenges and solutions.
+Figure 1: Our experiment design is detailed: we first conducted prompt engineering on the models (stage 1), and then, for better results, we selected a subset of models for supervised fine-tuning (stage 2).
+2 Related Work
+2.1 Hate Speech Detection
+Early studies established classifiers to detect harmful information using neural network models Zhang et al.
+(2018) or word embedding methods Kshirsagar et al.
+(2018).
+In recent years, models based on the Transformer architecture have demonstrated remarkable capabilities, prompting researchers to explore further.
+Rajput et al.
+(2021) conducted research on the ETHOS hate speech detection dataset, comparing classifiers’ performance in hate speech detection by replacing or integrating word embeddings (fastText, GloVe, or FT + GV) with BERT embeddings.
+Aluru et al.
+(2020) contrasted simple models (such as LASER embeddings with logistic regression) and BERT models in scenarios with scarce and abundant linguistic resources.
+Lin et al.
+(2024) generated explanations through multimodal debates between LLMs, enhancing the transparency and explainability of harmful meme detection.
+da Silva Oliveira et al.
+(2024) validated the effectiveness of ChatGPT in identifying harmful Spanish-language speech.
+2.2 Data Synthesis
+Traditional data synthesis methods have employed various techniques, ranging from synonym replacement to token-level manipulations, as exemplified by Wei and Zou (2019).
+While these methods provide some level of augmentation, they often lack contextual richness and diversity.
+The introduction of translation-based approaches (Fadaee et al., 2017) and masked language modeling (Kumar et al., 2020) improved semantic consistency, yet they still struggle to generate high-quality synthetic data necessary for complex tasks such as harmful content detection.
+Recent advancements in zero-shot data generation, such as ZEROGEN (Ye et al., 2022c), SuperGen (Meng et al., 2022), and PROGEN (Ye et al., 2022a), have explored methods for dataset synthesis without requiring extensive labeled examples.
+However, these frameworks often suffer from issues related to low information density and redundancy, limiting their effectiveness in generating diverse and contextually relevant samples.
 3 Methodology
-Our approach utilizes prompt engineering and fine-tuning techniques on open-source LLMs to enhance harmful data synthesis, thereby improving toxic content detection systems.
-The methodology comprises two key stages: prompt engineering and fine-tuning, as illustrated in Figure 1.
+Our methodology systematically evaluates open-source LLMs for harmful data synthesis through a two-stage approach: prompt engineering and supervised fine-tuning., as illustrated in Figure 1.
 3.1 Prompt Engineering
@@
 Our findings showed that relying solely on prompt engineering was inadequate for our objectives.
-This realization prompted the integration of a second stage—fine-tuning—to enhance the model’s ability to generate high-quality harmful data by retraining its weights with carefully curated datasets.
-2.2 Fine-Tuning
-Based on the results from prompt engineering, we hypothesized that fine-tuning would yield a more effective solution.
+This realization prompted the integration of a second stage—supervised fine-tuning—to enhance the model’s ability to generate high-quality harmful data by retraining its weights with carefully curated datasets.
+3.2 supervised fine-tuning
+Based on the results from prompt engineering, we hypothesized that supervised fine-tuning would yield a more effective solution.
 To achieve this, we employed LoRA methods Hu et al.
-(2022) for fine-tuning, updating the weights of $\mathcal{M}$ using the training dataset $\mathcal{D}_{\text{train}}$ to enhance the quality and diversity of $\mathcal{D}_{\text{harmful}}$.
-The fine-tuning objective, denoted as $\mathcal{L}$, minimizes the cross-entropy loss:
+(2022) for supervised fine-tuning, updating the weights of $\mathcal{M}$ using the training dataset $\mathcal{D}_{\text{train}}$ to enhance the quality and diversity of $\mathcal{D}_{\text{harmful}}$.
+The supervised fine-tuning objective, denoted as $\mathcal{L}$, minimizes the cross-entropy loss:
 $\mathcal{L}=-\sum_{i=1}^{N}\sum_{j=1}^{|V|}y_{ij}\log(\hat{y}_{ij})$
 where $y_{ij}$ represents the true label for the $j$-th vocabulary class of the $i$-th sample, $\hat{y}_{ij}$ is the predicted probability for the same, $|V|$ is the size of the vocabulary, and $N$ is the number of samples in the dataset.
-After fine-tuning, Mistral demonstrated the best performance compared to the baselines; the detailed experimental results are presented in Section 4.
+After supervised fine-tuning, Mistral demonstrated the best performance compared to the baselines; the detailed experimental results are presented in Section 5.
 3.3 Algorithm
@@
 14:return $\mathcal{D}_{\text{augmented}}$, $\mathcal{M}_{\text{down}}$
-Here, $\mathcal{D}_{\text{train}}$ represents the initial training dataset used for fine-tuning, $\mathcal{D}_{\text{harmful}}$ consists of unique harmful data samples generated by the fine-tuned LLM, and $\mathcal{D}_{\text{augmented}}$ is the final dataset combining $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{harmful}}$ for downstream model training.
+Here, $\mathcal{D}_{\text{train}}$ represents the initial training dataset used for supervised fine-tuning, $\mathcal{D}_{\text{harmful}}$ consists of unique harmful data samples generated by the fine-tuned LLM, and $\mathcal{D}_{\text{augmented}}$ is the final dataset combining $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{harmful}}$ for downstream model training.
 Duplicates in $\mathcal{D}_{\text{harmful}}$ are removed.
@@
 Mix_GPT (Epoch 5) 0.667 0.734 0.802 0.852 0.702 0.749 0.788 0.794 - -
-Table 3: Performance Comparison of Fine-Tuned Models on Different Datasets, best results are indicate in bold.
+Table 3: Performance Comparison of Fine-Tuned Models on Different Datasets, best results are indicated in bold.
 4.2 Models
-Stage 1: Prompt Engineering
-n the first stage of our study, we employed prompt engineering to systematically evaluate the performance of six open-source LLMs: Mistral Jiang et al.
+Prompt Engineering
+In this stage of our study, we employed prompt engineering to systematically evaluate the performance of six open-source LLMs: Mistral Jiang et al.
 (2023), LLaMA2 Touvron et al.
@@
 (2023), and Gemma Team et al.
+(2024) and our baseline: GPT4 model Brown et al.
 (2020).
 Detailed results and analysis for Stage 1 are provided in Section 5.1.
-Based on the performance evaluation, Mistral and Vicuna were identified as the top-performing models in the prompt engineering stage, and were subsequently selected for the fine-tuning stage.
-Stage 2: Fine-Tuning
-In the second stage, we fine-tuned the Mistral and Vicuna models using the LoRA method Hu et al.
+Based on the performance evaluation, Mistral and Vicuna were identified as the top-performing models in the prompt engineering stage, and were subsequently selected for the supervised fine-tuning stage.
+Supervised Fine-tuning
+In supervised fine-tuning, we fine-tuned the Mistral Jiang et al.
+(2023) and Vicuna Chiang et al.
+(2023) models using the LoRA method Hu et al.
 (2022).
-After comparing latency and cost we decided to use a smaller model Mixtral for further exploration.
+After comparing latency and cost, we experimented and reported the better performance model Mixtral for further exploration.
 Detailed results and analysis for Stage 2 are provided in Section 5.2.
 4.3 Implementation
-The fine-tuning process was conducted on an Azure cluster equipped with 4 * NVIDIA A100 GPUs, leveraging the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework Zheng et al.
+The supervised fine-tuning process was conducted on an Azure cluster equipped with 4 * NVIDIA A100 GPUs, leveraging the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework Zheng et al.
 (2024).
@@
 Detailed training parameters, including batch size, learning rate, and optimizer settings, are provided in the Appendix B.
-The results of this fine-tuning process, along with performance metrics and qualitative analyses, are presented in Section 4.
+The results of this supervised fine-tuning process, along with performance metrics and qualitative analyses, are presented in Section 5.
 Downstream Model Evaluation
-In our production environment, meeting strict inference time requirements is critical.
-Among the evaluated models, the Multilayer Perceptron (MLP) Gardner and Dorling (1998) stood out for its ability to fulfill these demands.
-For downstream model evaluation, in this paper we utilized a MLP classifier with two hidden layers, each comprising 600 neurons.
-4 Result
-This section presents the results of our study, showcasing the effectiveness of both prompt engineering and fine-tuning strategies in optimizing open-source LLMs for harmful data synthesis.
-Our evaluation employed a two-stage approach: Stage 1 focused on prompt engineering to assess the initial performance of various models, while Stage 2 involved fine-tuning the selected models to further enhance data quality and diversity.
+We utilized a Multilayer Perceptron (MLP) Gardner and Dorling (1998) for downstream model evaluation.
+The MLP classifier consisted of two hidden layers, each comprising 600 neurons, with ReLU activation functions and a softmax output layer for classification.
+The model was trained using the Adam optimizer.
+We conducted multiple training runs with different random seeds to ensure stability and robustness in evaluation.
+Given its simple structure and strong interpretability, MLP provides a clear and reliable baseline for assessing the impact of synthetic data on model performance across various datasets and experimental conditions.
+5 Results and Analysis
+We present the results of our study, structured around the three research questions.
+Our two-stage evaluation—prompt engineering and supervised fine-tuning—allowed us to assess the effectiveness of open-source LLMs in synthetic toxic data generation, examine the limitations of prompt-based methods, and analyze the impact of different generation strategies on model robustness.
 5.1 Stage 1: Prompt Engineering
-The first stage of our evaluation focused on assessing the performance of various Open Source LLMs for data synthesis using prompt engineering.
-The primary goal was to explore the quality and effectiveness of existing LLMs in generating augmented data for political content.
-Table 2 presents the performance comparison of the six different models evaluated in Stage 1.
-The success rates for political and hate content are reported, along with the results of human evaluation.
-The success rate columns in Table 2 represent the percentage of valid positive samples generated by each model for political and hate content.
-The values indicate the proportion of samples meeting the minimum requirements (60% of the total samples generated).
-The human evaluation columns show whether the generated data was subject to human evaluation.
-A green checkmark () signifies that human annotators evaluated the quality of the generated data and found a high Fleiss’ kappa score, indicating strong inter-rater agreement.
-This also means that the generated samples met the minimum quality criteria.
-Based on the evaluation, Mistral and Vicuna demonstrated the best performance with prompt engineering.
-These two models were selected for the fine-tuning stage due to their superior results in generating high-quality and diverse data entries.
+The first stage of our evaluation assessed how well open-source LLMs could generate synthetic data using prompt engineering.
+The primary objective was to evaluate their ability to generate high-quality augmented data, particularly for political and hate speech content.
+Table 2 presents a performance comparison of six open-source models, including their success rates for political and hate content as well as human evaluation results.
+The success rate columns in Table 2 indicate the percentage of valid positive samples generated by each model, where a minimum threshold of 60% of the total samples was required.
+The human evaluation columns confirm whether the generated data underwent human annotation.
+A green checkmark () signifies that annotators validated the data quality, achieving a high Fleiss’ kappa score, which indicates strong inter-rater agreement.
+5.2 Stage 2: Supervised Fine-tuning
+In the second stage, models were supervised fine-tuned on individual datasets and evaluated accordingly, with results detailed in Table 4.
+Inspired by Hui et al.
+(2024), Yu et al.
+(2024), and Qi et al.
+(2023), we then explored data mixing, incorporating multiple datasets for training.
+Table 3 presents results from this multi-dataset fine-tuning strategy.
+Further analysis, including an ablation study on different design choices, is provided in Figure  2 and Appendix D.
+The "Mixture" versions in Table 3 represent models trained on a blend of hate, violence, and sex-related datasets.
+The "Mix_GPT" versions extend this approach by integrating GPT-generated positive samples.
+Across all fine-tuning experiments, we maintained a balanced dataset of 3000 positive and 3000 negative samples, formatted using the Alpaca instruction template.
 Dataset Version F1 Score Accuracy
@@
 Sex_epoch1 0.587 0.702
-Table 4: Performance Comparison of Fine-Tuned Models on Hate and Sex, best result are indicate in bold.
-4.2 Stage 2: Fine-tuning
-In the second stage, we began by fine-tuning the models exclusively on individual datasets and evaluating their performance on the same datasets.
-The results of this approach are detailed in Table 4.
-Inspired by Hui et al.
-(2024), Yu et al.
-(2024), and Qi et al.
-(2023), we subsequently adopted a data mixing strategy that combined multiple datasets for fine-tuning.
-The results of this multi-dataset fine-tuning are presented in Table 3.
-Additionally, we conducted an ablation study to examine the impact of different design choices, detailed in Appendix D.
-The "Mixture" versions represent fine-tuning on a mix of hate, violence, and sex datasets, while the "Mix_GPT" versions incorporate GPT-generated positive samples into the mixed data.
-For all fine-tuning experiments, we used 3000 positive and 3000 negative samples, following the Alpaca instruction template.
-Our findings reveal that fine-tuning significantly enhances the performance of LLMs in generating harmful data.
-The data mixing strategy helps in mitigating issues like data duplication and overfitting.
-The fine-tuned models exhibited higher F1 scores and accuracy across different datasets, demonstrating their robustness and effectiveness in real-world applications.
-One surprising observation was the notable improvement in model performance when using mixed data, suggesting that diverse training data enables better generalization and more reliable outputs.
-The Mixture models trained for Epoch 3 and Epoch 5 showed particularly strong results, emphasizing the importance of sufficient training epochs for optimal performance.
+Table 4: Performance Comparison of Fine-Tuned Models on Hate and Sex, best results are indicated in bold..
+Figure 2: Ablation Study on Data mixing
+5.3 RQ1: How effectively can open-source LLMs generate synthetic toxic data compared to proprietary models?
+We compared six open-source LLMs against a GPT-4 baseline using prompt-based generation.
+Mistral and Vicuna consistently achieved the highest success rates and human evaluation scores among open models (Table 2).
+However, even these models struggled to match GPT-4 in data diversity and nuance, often producing repetitive or formulaic outputs.
+Fine-tuning significantly reduced this gap.
+A fine-tuned Mistral model demonstrated improved F1 scores and accuracy across multiple datasets (Table 4).
+Although GPT-4 remained the strongest performer, fine-tuned Mistral approached its performance, demonstrating that open-source LLMs, when fine-tuned, can serve as viable alternatives for synthetic data generation.
+5.4 RQ2: What limitations arise from using prompt-based synthetic data generation for hate speech detection, and how can fine-tuning improve performance?
+Prompt-based generation exhibited several critical limitations.
+Safety alignments in models such as Gemma and Bloom led to frequent refusals to generate toxic content, reducing dataset completeness.
+Additionally, generated samples often lacked diversity, limiting their utility for training robust detection models.
+Another issue was over-simplicity, as models struggled to produce nuanced toxicity, instead generating overly blunt or obvious harmful statements.
+Fine-tuning alleviated these challenges.
+Fine-tuned models demonstrated improved data diversity, generating more varied and contextually rich toxic content (Table 3).
+They also exhibited higher data reliability, reducing hallucination rates and enhancing dataset quality.
+Furthermore, fine-tuning enabled models to generate more subtle and realistic toxic content, making them more applicable to real-world hate speech detection.
+However, fine-tuning also introduced challenges such as data duplication and overfitting, particularly in later training epochs (Table 4).
+5.5 RQ3: How do different synthetic data generation strategies impact the robustness of toxic content detection models?
+We analyzed how different synthetic data generation strategies influenced classification performance.
+Fine-tuned models consistently outperformed those relying solely on prompt-generated data, achieving higher F1 scores and accuracy across multiple datasets (Table 4).
+Models trained on blended datasets containing multiple categories of toxic content (e.g., hate speech, violence, and sexual content) demonstrated stronger generalization compared to models trained on a single category.
+However, extended fine-tuning introduced overfitting risks.
+Performance improvements plateaued beyond epoch 3, and some models experienced performance degradation at epoch 5 (Table 4), likely due to memorization effects.
+Ablation studies (Appendix D) confirmed that strategic fine-tuning with mixed data provides the best balance between data quality and generalization, underscoring the importance of careful dataset curation and tuning strategies.
 5.6 Error Analysis
-We conducted an error analysis on the fine-tuned models to better understand the nature of their mistakes and identify areas for improvement.
+We conducted an error analysis on the fine-tuned models to better understand their mistakes and identify areas for improvement.
 This analysis focuses on two key aspects: (1) generation errors and (2) classification errors.
+Generation errors refer to cases where the model produces responses that are misleading, incomplete, or inconsistent with the input data.
+Classification errors occur when the model incorrectly labels content, either failing to detect harmful content (false negatives) or misclassifying benign content as harmful (false positives).
 Generation Errors.
-The data generation process was evaluated for repetition rates and the nature of outputs across different epochs and model versions for the hate and sex datasets.
-For example Mixtral_Mixture3 models were effective in generating targeted harmful data but occasionally produced random string like " You are 32\$\$2\$(2djs…" , which could skew downstream training.
-Additionally, repetitive outputs across epochs indicated potential overfitting in certain scenarios, suggesting the need for better strategies to ensure diversity and control the tone of generated content.
+The data generation process was evaluated for repetition rates and output consistency across different epochs and model versions.
+For example, the Mixtral_Mixture3 model effectively generated targeted harmful data but sometimes produced malformed text, such as random character sequences or redundant repetitions.
+Below, we present five representative cases of generation errors.
+Examples of Generation Errors 1.
+Random String Noise: "You are 32\$\$2\$(2djs..." (Model unexpectedly generates meaningless characters, likely due to tokenization errors.) 2.
+Repetitive Output: "You are a fool.
+You are a fool.
+You are a fool." (Overfitting leads to excessive repetition across different epochs.) 3.
+Overfitting Issue: "Get out of my sight." (Identical phrasing appears in multiple generations, suggesting memorization.) 4.
+Unintended Politeness: "Sir, I would like to insult you, but I must remain polite." (LLM exhibits misalignment between intent and expression.) 5.
+Formatting Glitch: "!!!##You are nothing but##sh*t#(!!" (Generated output includes unnecessary symbols or special characters.)
 Classification Errors.
-We examined classification errors in the downstream MLP classifier trained on augmented data.
-Examples of missed classifications include statements like, "You are a disabled person," which might have been overlooked due to its less explicit tone compared to direct threats.
-Similarly, a statement like, "I’m sorry you’re deaf," may have been misclassified as non-harmful due to its seemingly apologetic phrasing.
-These cases highlight the difficulty of identifying subtle harmful content and underscore the need for classifiers capable of handling nuanced language.
+We analyzed misclassified examples in the downstream MLP classifier trained on augmented data.
+The most challenging errors involve implicit harm, sarcasm, and contextual misinterpretation.
+Below, we present three representative cases with detailed explanations.
+Examples of Classification Errors 1.
+Sarcastic Threat (False Negative): "Oh wow, you’re such a genius.
+Maybe I should come over and personally teach you some ‘respect’—don’t worry, it won’t hurt...
+much." (This statement uses heavy sarcasm ("such a genius") while masking a direct physical threat ("personally teach you some ‘respect’").
+The phrase "don’t worry, it won’t hurt… much" adds a sinister undertone, but due to the sarcastic phrasing, the classifier failed to recognize it as harmful.) 2.
+Context Misinterpretation (False Positive): "That’s a killer move in chess!" (The model incorrectly flagged this as harmful due to the word "killer," failing to recognize the benign context in a strategic game discussion.)
+6 Online Deployment
+The fine-tuned models developed in this study have been successfully deployed in real-world applications by a leading cloud provider, enhancing automated content moderation workflows.
+These models contribute to real-time content filtering, reducing manual intervention while improving accuracy and efficiency.
+Deployment results indicate that fine-tuned open-source LLMs can effectively identify and flag harmful content, demonstrating strong performance in diverse online environments.
+Key deployment challenges include latency, scalability, and adaptability to evolving harmful content patterns.
+Our models were optimized to operate within industry constraints, ensuring rapid inference while maintaining high detection accuracy.
+Additionally, continuous monitoring and retraining mechanisms have been implemented to adapt to emerging harmful language trends, reinforcing the robustness of our approach.
+7 Conclusion
+Our findings underscore both the potential and challenges of using open-source LLMs for toxic content synthesis.
+The two-stage evaluation approach highlighted key trade-offs between prompt engineering and supervised fine-tuning.
+While prompt engineering offers a rapid and lightweight approach, its effectiveness is limited by the inherent safety mechanisms within LLMs, which restrict harmful content generation.
+Fine-tuning, on the other hand, significantly improves diversity and realism but introduces risks such as overfitting and data redundancy.
+A notable challenge is the subtlety of harmful content.
+Fine-tuned models struggled with detecting nuanced harmful language, such as implicit bias and sarcasm.
+Additionally, repetition and overfitting were observed in later training epochs, indicating that fine-tuning requires careful calibration to balance specificity and generalization.
+Our data mixing strategy, which incorporated multiple categories of harmful content, demonstrated benefits in improving generalization.
+However, maintaining a balance between specificity and coverage remains a key challenge, especially for mixed-content datasets.
+Further research is required to refine fine-tuning methodologies, ensuring that generated content remains relevant, diverse, and applicable across different domains.
 8 Discussion
-The results of this study highlight the complexities and challenges associated with generating and detecting harmful content using LLMs.
-Through a two-stage evaluation process, we examined the performance of various open-source LLMs and assessed the impact of fine-tuning on mixed datasets.
-Several key insights emerged: One of the primary challenges observed was the subtlety of harmful content.
-The fine-tuned models often failed to detect subtle harmful statements that were not overtly aggressive or violent.
-This limitation underscores the need for more sophisticated detection mechanisms capable of understanding nuanced language and contextual subtleties.
-Repetition and overfitting were also significant issues, particularly in the later epochs of the fine-tuned models.
-High repetition rates suggest that the models may be memorizing specific phrases rather than generalizing effectively across different harmful contexts.
-This overfitting reduces the diversity and overall effectiveness of harmful content generation.
-The data mixing strategy, which combined datasets of hate, violence, and sex with GPT-generated positive samples, showed promising results in improving both precision and recall.
-However, the models continued to struggle with maintaining a balance between specificity and generalization, especially when dealing with mixed types of harmful content.
-6 Online Deployment
-The models developed in this study have been successfully deployed in a real-world application by a leading cloud provider, offering valuable insights and automations for a range of tasks.
-Specifically, these models are used to enhance content moderation workflows, enabling faster and more accurate detection of harmful content in real-time scenarios.
-Their deployment has demonstrated the practicality of fine-tuned open-source LLMs in addressing challenges such as latency, scalability, and adaptability to diverse content types.
-7 Conclusion
-This study showcases the potential of open-source LLMs for harmful data synthesis and their practical applications in toxic content detection systems.
-By evaluating prompt engineering and fine-tuning approaches, we identified the strengths and limitations of various models in generating diverse, high-quality harmful data.
-Our findings underline the importance of balancing data quality and diversity with computational efficiency to develop scalable and cost-effective content moderation solutions.
-Fine-tuning emerged as a key factor in enhancing model performance.
-Through our data mixing strategy, we successfully addressed challenges like data duplication and overfitting, improving model generalization and robustness.
-However, challenges such as detecting subtle harmful content and balancing specificity with generalization remain open issues that warrant further research.
-Future work should explore innovative techniques to address these challenges, such as incorporating contextual embeddings, leveraging multimodal data sources, and experimenting with advanced fine-tuning methods.
-Additionally, our findings highlight the potential of fine-tuned open-source LLMs to enable cost-effective solutions for industrial applications, making harmful data synthesis and detection tools more accessible to smaller organizations.
+Future research should focus on improving adaptability, contextual awareness, and fairness in toxic content detection.
+Dynamic adaptation to emerging toxic content is crucial, as harmful expressions evolve rapidly.
+Self-supervised continual learning can help models detect new toxic patterns without frequent retraining.
+Improving contextual awareness is another key challenge, as toxic language often relies on implicit meaning.
+Integrating external knowledge graphs and multimodal signals (text, images, audio) can enhance models’ ability to understand nuanced toxicity.
+Reducing overfitting and enhancing diversity is essential to prevent models from becoming too specialized.
+Techniques like contrastive learning, reinforcement learning, and curriculum learning can help balance specificity and generalization.
+Interactive human-in-the-loop systems can further improve accuracy and scalability by incorporating expert feedback into real-time model updates.
+Lastly, ethical alignment and bias mitigation require fairness-aware training methodologies and adversarial testing frameworks to minimize unintended biases in synthetic toxic data generation.
+9 Responsible AI Statement
+Our study acknowledges the ethical considerations and potential risks associated with generating synthetic harmful content.
+While our research aims to improve automated content moderation, we recognize that models capable of generating toxic data could be misused.
+To mitigate these risks, we adopt strict safety measures, including controlled access to fine-tuned models, comprehensive dataset curation to prevent unnecessary amplification of harmful language, and continuous monitoring for unintended biases.
+Additionally, we emphasize the need for transparent documentation and responsible deployment practices to ensure that synthetic data generation aligns with ethical AI principles.
+Future work should further explore fairness-aware training methodologies, adversarial testing frameworks, and regulatory oversight to minimize potential harms while maximizing societal benefits.
 10 Limitations
 While our study provides valuable insights into the performance of fine-tuned LLMs for harmful content generation and detection, several limitations must be acknowledged.
-One of the primary limitations is the diversity and representation of the datasets used for fine-tuning.
+One of the primary limitations is the diversity and representation of the datasets used for supervised fine-tuning.
 Although we utilized mixed datasets combining hate, sex, and violence categories, the scope remains limited to specific types of harmful content.
@@
 Ensuring the responsible use of LLMs in generating harmful content is crucial, and future research should focus on developing more sophisticated ethical guidelines and safety measures.
-The fine-tuning process in this study relied on significant computational resources, including multiple A100 GPUs.
+The supervised fine-tuning process in this study relied on significant computational resources, including multiple A100 GPUs.
 This requirement may not be feasible for all researchers or applications, particularly those with limited access to high-performance computing resources.
-Exploring more efficient fine-tuning techniques and leveraging less resource-intensive models would be beneficial for broader accessibility and scalability.
+Exploring more efficient supervised fine-tuning techniques and leveraging less resource-intensive models would be beneficial for broader accessibility and scalability.
 Ethics Statement
@@
 Appendix B Detailed Training Parameters
-The fine-tuning process was conducted on an Azure cluster equipped with 4 NVIDIA A100 GPUs, utilizing LLaMA-Factory Zheng et al.
+The supervised fine-tuning process was conducted on an Azure cluster equipped with 4 NVIDIA A100 GPUs, utilizing LLaMA-Factory Zheng et al.
 (2024).
@@
 Appendix D Ablation Study
-In this section, we explore the effects of using mixed data for fine-tuning models and compare the results using different versions of prompts (v1, v2, v3).
+In this section, we explore the effects of using mixed data for supervised fine-tuning models and compare the results using different versions of prompts (v1, v2, v3).
 For simplification, the effects of different prompt versions are considered the same across all experiments.
@@
 Comparison with Single-category Data: Models fine-tuned on single-category data (e.g., hate_v1, sex_v1) show high precision but lower recall compared to those fine-tuned on mixed data.
-This indicates that single-category data fine-tuning might lead to more conservative models that are less likely to generalize to other types of harmful content.
+This indicates that single-category data supervised fine-tuning might lead to more conservative models that are less likely to generalize to other types of harmful content.
 Prompt Versions: The experiments using different prompt versions (v1, v2, v3) show similar performance metrics, indicating that the choice of prompt version has a minimal impact on the overall performance of the fine-tuned models.
-The ablation study demonstrates that fine-tuning models with mixed datasets (hate_multi) can significantly improve recall and F1 scores, enhancing the models’ ability to detect various types of harmful content.
+The ablation study demonstrates that supervised fine-tuning models with mixed datasets (hate_multi) can significantly improve recall and F1 scores, enhancing the models’ ability to detect various types of harmful content.
 However, there is a trade-off with precision, requiring careful consideration in applications where false positives are costly.